### v1

# Bias Amplification: Language Models as Increasingly Biased Media

Abstract
As Large Language Models (LLMs) become increasingly integrated into various facets of society, a significant portion of online text consequently become synthetic. This raises concerns about bias amplification, a phenomenon where models trained on synthetic data amplify the pre-existing biases over successive training iterations. Previous literature seldom discusses bias amplification as an independent issue from model collapse. In this work, we address the gap in understanding the bias amplification of LLMs with four main contributions. Firstly, we propose a theoretical framework, defining the necessary and sufficient conditions for its occurrence, and emphasizing that it occurs independently of model collapse. Using statistical simulations with weighted maximum likelihood estimation, we demonstrate the framework and show how bias amplification arises without the sampling and functional form issues that typically drive model collapse. Secondly, we conduct experiments with GPT-2 to empirically demonstrate bias amplification, specifically examining open-ended generational political bias with a benchmark we developed. We observe that GPT-2 exhibits a right-leaning bias in sentence continuation tasks and that the bias progressively increases with iterative fine-tuning on synthetic data generated by previous iterations. Thirdly, we explore three potential mitigation strategies: Overfitting, Preservation, and Accumulation. We find that both Preservation and Accumulation effectively mitigate bias amplification and model collapse. Finally, using novel mechanistic interpretation techniques, we demonstrate that in the GPT-2 experiments, bias amplification and model collapse are driven by distinct sets of neurons, which aligns with our theoretical framework.
Bias Amplification: Language Models as Increasingly Biased Media
Ze Wang1,2, Zekun Wu1,2, Jeremy Zhang3, Navya Jain 2 Xin Guan 1, Adriano Koshiyama 1 1Holistic AI, 2University College London, 3Emory University, Correspondence: ze.wang.19@ucl.ac.uk, zekun.wu@holisticai.com
1 Introduction
Large language models (LLMs) are trained on vast amounts of text scraped from the internet, which plays a crucial role in improving their capabilities, whether through emergent abilities (Wei et al., 2022) or scaling laws (Kaplan et al., 2020). However, as LLMs become more widely integrated into human society—for example, in content creation and summarization in media, academia, and business (Maslej et al., 2024)—concerns are mounting that a significant portion of online text in the future may be generated, either entirely or partially, by LLMs (Peña-Fernández et al., 2023; Porlezza and Ferri, 2022; Nishal and Diakopoulos, 2024). This highlights a significant and underexplored risk: bias amplification, referring to the degradation of fairness within LLMs over self-consuming training loops (Mehrabi et al., 2022; Taori and Hashimoto, 2022), where models progressively amplify the pre-existing biases. This concern initially arises from the tendency of LLMs to learn from biased datasets. For example, Parrish et al. (2022); Wang et al. (2024); Bender et al. (2021) show that LLMs absorb inherent stereotypes, such as racial and gender discrimination, embedded in human-generated text. Additionally, Haller et al. (2023); Rettenberger et al. (2024a) demonstrated that LLMs can be aligned with specific political ideologies by fine-tuning them on biased datasets. Moreover, Wyllie et al. (2024) shows that classifiers trained on synthetic data increasingly favor a particular class label over successive generations, while the shrinking diversity observed by Alemohammad et al. (2023); Hamilton (2024) suggests a risk that certain demographic groups may become progressively underrepresented in the outputs of LLMs.
The amplification of biases has profound societal implications. It can lead to the perpetuation of stereotypes, reinforcement of social inequalities, and the marginalization of underrepresented groups. In the context of political bias, this can influence public opinion, skew democratic processes, and exacerbate polarization. Understanding and mitigating bias amplification is therefore crucial to ensure that LLMs contribute positively to society and do not inadvertently cause harm. Nevertheless, despite the literature on bias amplification in discriminative models, there is a notable lack of comprehensive frameworks and empirical studies specifically addressing bias amplification for LLMs, as shown in Section 2. In contrast to discriminative models, where bias amplification can be attributed to overfitting the dominant features in the training dataset, its occurrence in LLMs may stem from a more nuanced issue. As we will show in Sections 3 and 5, even when the training data is unbiased, the model can still amplify its pre-existing biases during the training process.
In this paper, we seek to fill this research gap by proposing a theoretical framework that establishes and explains the causes of bias amplification in LLMs. Additionally, we perform both statistical simulation and LLM experiments to demonstrate and exemplify the theorem. In summary, our main contributions are as follows:
1. Theoretical Framework: We establish the necessary and sufficient conditions for bias amplification (i.e. Theorem 1). The theorem aids in understanding the cause of bias amplification and in distinguishing between bias amplification and model collapse. We conduct statistical simulations using weighted maximum likelihood estimation to illustrate the theorem (see Section 3.2).
2. Benchmarking Tool: We trained a highly accurate classifier capable of detecting political leaning in long-text content. With this classifier, we offer a benchmark for evaluating political bias in LLMs through open-generation tasks, filling a gap not covered in current bias studies (see Sections 2 and 4.3).
3. Empirical Demonstration and Mitigation Strategies: We demonstrate bias amplification in terms of political bias in GPT-2 using our benchmarking tool: the model exhibits a right-leaning bias in sentence continuation tasks and becomes increasingly right-leaning over successive generations (see Section 5.1). Additionally, we conducted experiments with three potential mitigation strategies, i.e. Overfitting, Preservation, and Accumulation, comparing their effectiveness, and found that some are surprisingly effective (see Section 5.3).
4. Mechanistic Interpretation: Building on our framework, we propose an innovative mechanistic interpretation pipeline that identifies two distinct sets of neurons responsible for bias amplification and model collapse during iterative fine-tuning experiments with GPT-2. We found minimal overlap between these two sets, supporting our theorem that bias amplification can occur independently of model collapse (see Section 5.4).
The rest of the paper is organized as follows: We first discuss related work in Section 2. In Section 3, we present our theoretical framework for bias amplification and conduct a statistical simulation. Section 4 describes the experimental setup for our LLM experiments, including data preparation and model fine-tuning. In Section 5, we present our empirical findings on bias amplification and the effectiveness of mitigation strategies. Finally, in Section 6, we discuss the implications of our work, concluding with limitations in Section 7.
2 Background and Related Work
Bias Amplification. Bias amplification in machine learning models has been seriously studied in the context of visual recognition tasks. For example, Zhao et al. (2017) found that Conditional Random Fields can exacerbate social biases present in the training data. This study proposed an in-process mitigation approach, employing Lagrangian Relaxation to enforce constraints that ensure the model’s bias performance remains closely aligned with the biases in the training data. Following this, Mehrabi et al. (2022) proposed the concept of bias amplification in feedback loops, where biased models not only amplify the bias present in their training data but also interact with the world in ways that generate more biased data for future models. Xu et al. (2023); Zhou et al. (2024) examined bias amplification in recommendation models, showing how these models reinforce their understanding of mainstream user preferences from training data, leading to an overrepresentation of such preferences in historical data and neglecting rarely exposed items—similar to the concept of sampling error discussed in (Shumailov et al., 2024). Additionally, Wyllie et al. (2024); Taori and Hashimoto (2022) demonstrated that classifiers trained on synthetic data increasingly favor specific class labels over successive generations. Likewise, Ferbach et al. (2024); Chen et al. (2024) observed bias amplification in generative models such as Stable Diffusion, characterized by the overrepresentation of features from the training dataset. Moreover, we conducted a comprehensive review of the literature on model collapse which we put in Appendix L.
Political Biases. LLMs become a new form of "media" that people rely on for global news (Maslej et al., 2024), attracting significant research attention to political biases. Rettenberger et al. (2024b); Shumailov et al. (2024); Feng et al. (2024) explored the bias through voting simulations within the spectrum of German political parties, consistently finding a left-leaning bias in models like GPT-3 and Llama3-70B. Similarly, for the U.S. political landscape, Rotaru et al. (2024); Motoki et al. (2024) identified a noticeable left-leaning bias in ChatGPT and Gemini when tasked with rating news content, evaluating sources, or responding to political questionnaires.
3 Theoretical Framework
In this section, we formalize the conditions under which bias amplification occurs during recursive training on synthetic data. This framework is designed to intuitively illustrate the primary causes of bias amplification. Next, we demonstrate bias amplification and the theorem using Weighted Maximum Likelihood Estimation (WMLE).
3.1 The Conditions for Bias Amplification
Bias amplification can arise from two key factors. The first is referred to as bias projection. It arises when the bias projection coefficient (explained below) is negative, indicating that the gradient update during training will fully align with the existing bias in the model parameters, thus amplifying it. Consider a fine-tuning process where the pre-trained model parameters $\theta_{t}$ can be decomposed into unbiased and biased components:
$\theta_{t}=\theta_{t,\text{unbiased}}+\theta_{t,\text{biased}}.$
During gradient-based optimization, the update rule is:
$\theta_{t+1}=\theta_{t}-\eta\nabla_{\theta}\mathcal{L}_{\text{ft}}(\theta_{t}),$
where $\eta$ is the learning rate, and $\mathcal{L}_{\text{ft}}$ denotes the fine-tuning loss function. Substituting the decomposition of $\theta_{t}$ and taking the projection, we have:
$\theta_{t+1}=\theta_{t,\text{unbiased}}+\theta_{t,\text{biased}}-\eta\left(\frac{\mathbf{\theta_{t,\text{biased}}}}{\|\mathbf{\theta_{t,\text{biased}}}\|}\right)c_{t}$
where $c_{t}$ is the bias projection coefficient, measuring the projection of the gradient onto the normalized biased component of the parameters:
$c_{t}=\left(\frac{\theta_{t,\text{biased}}}{\|\theta_{t,\text{biased}}\|}\right)^{\top}\nabla_{\theta}\mathcal{L}_{\text{ft}}(\theta_{t}).$ (1)
If $c_{t}<0$, the gradient update will reinforce the biased component, leading to bias amplification, i.e. $\Delta|\theta_{\text{biased}}|>0$. This occurs because the gradient descent step moves the parameters further in the direction of the existing bias. Intuitively, the bias in the output is influenced by specific neurons, and increasing the weights of these neurons best aligns with the training objective for the given dataset. This does not require the dataset itself to be biased, but rather that it contains features, aside from bias, that align more closely with the functionalities of biased neurons relative to the unbiased neurons considering the loss function.
The second factor is referred to as constraint deficiency. It occurs when the loss or objective function lacks explicit constraints or regularization to counteract bias projection. As a result, the optimization process does not prevent the reinforcement of bias when bias projection is present. To prevent bias amplification theoretically, the loss function can incorporate de-biasing constraints. Consider a modified fine-tuning objective: $\mathcal{L}_{\text{new}}(\theta)=\mathcal{L}_{\text{original}}(\theta)-\lambda\sum_{i=1}^{n}\max(0,c^{i})^{2}+\gamma\sum_{i=1}^{n}\min(0,c^{i})^{2}$, where $c^{i}$ represents the bias projection coefficient for a specific dimension. $\lambda,\gamma\geq 0$ are the Lagrangian multipliers for the two constraints:
(1) $\sum_{i=1}^{n}\max(0,c^{i})^{2}\geq 0$,   i.e. encouraging positive components.
(2) $\sum_{i=1}^{n}\min(0,c^{i})^{2}\leq 0$,  i.e. eliminating negative components.
The absence of such constraints allows bias amplification when bias projection is present. Based on these findings, we establish the following:
Theorem 1 (Condition for Bias Amplification).
Bias amplification, a self-reinforcing process where a model trained on synthetic data amplifies the pre-existing biases from previous training, leading to an increased bias in the data it generates for future training, occurs if both bias projection and constraint deficiency are present. It occurs only if there is bias projection in at least one parameter dimension and constraint deficiency is present.
Alongside understanding the causes of bias amplification, this also implies that bias amplification can occur even without model collapse. Specifically, statistical approximation and functional expressivity errors [^1] We have not discuss functional approximation error here, as it is a general concept that encompasses any structural setting, from learning to objective function, that may lead to suboptimal training outcomes, making it too general for this discussion.—common causes of model collapse (Shumailov et al., 2024)—are unrelated with bias projection and constraint deficiency. This is further supported by our experiments in Sections 3.2 and 5.
3.2 Weighted Maximum Likelihood Estimation (WMLE)
To demonstrate bias amplification in a controlled setting, we consider a statistical estimation cycle using WMLE, devoid of statistical approximation and functional expressivity errors. The mathematical formulation of WMLE is provided in Appendix A. Specifically, we generate a "biased" pre-training dataset $\mathcal{D}_{\text{pre}}$ consisting of $N=10,000$ samples from a "biased" Beta$(\alpha=3,\beta=2)$ distribution, with a mean of $x=0.6$. Maximum likelihood estimation (MLE) is used to estimate the probability density function (pdf) of this beta distribution, resulting in the ’biased’ pre-trained pdf (referred to as $f_{\text{pre}}$). Our goal in the estimation cycle is to estimate the pdf of the true distribution, Beta$(\alpha=2,\beta=2)$, with a mean of $x=0.5$. We generate $N$ samples from the true distribution, referred to as $D_{\text{real}}$, serving as data for the initial round of training. In the first round, we estimate the true pdf using WMLE with $D_{\text{real}}$ and weights derived from $f_{\text{pre}}$, simulating the impact of the biased pre-trained model during the current training. The newly estimated pdf is referred to as $f_{\text{Generation 0}}$. Then, we generate the synthetic data $D_{\text{Generation 0}}$ of size N using $f_{\text{Generation 0}}$. In the second round, we use WMLE with $D_{\text{Generation 0}}$ and weights derived from $f_{\text{Generation 0}}$ to obtain $f_{\text{Generation 1}}$. We then repeat this process, recursively estimating $f_{\text{Generation 2}}$ through $f_{\text{Generation 10}}$, simulating an iterative training process on synthetic data.
Results:
As depicted in Figure 1, the estimated distributions progressively shift towards the mean of the biased Beta$(\alpha=3,\beta=2)$ distribution, at $x=0.6$, becoming increasingly peaked with each generation. This behavior exemplifies bias amplification. The distortion arises from two sources: bias projection, where the estimation process emphasizes data regions where the ’biased’ pre-trained distribution assigns higher probability; and constraint deficiency, where WMLE lacks mechanisms to counteract the bias introduced by the weights, allowing the bias to be amplified over generations. For comparison, Figure 2 shows the results using standard MLE without weighting. The estimated distributions remain consistent across generations, accurately capturing the true Beta(2,2)22(2,2)( 2 , 2 ) distribution. This indicates that without the influence of biased pre-trained pdf and the weighted learning process, bias amplification does not occur.
Figure 1: Weighted Maximum Likelihood Estimation (WMLE) over 10 generations.
Figure 2: Standard Maximum Likelihood Estimation (MLE) over 10 generations, starting from the true distribution Beta(2,2)22(2,2)( 2 , 2 ) without weighting.
4 Experiment Design
In this section, we present and explain in detail the experiments designed to investigate the bias amplification of LLMs (see results in Section 5). Specifically, we analyze the sequential and synthetic fine-tuning of GPT-2. While similar experiments could be conducted with larger models such as GPT-4 or Claude in non-fine-tuning contexts, doing so would likely incur significant environmental costs. The focus of this study is on political bias within the US political spectrum, particularly in the context of sentence continuation tasks. This is important because LLMs are increasingly becoming a source of information for global news consumption Maslej et al. (2024); Peña-Fernández et al. (2023); Porlezza and Ferri (2022), and traditional news outlets, such as the Associated Press, are beginning to incorporate LLMs to automate content generation from structured data (The Associated Press, 2024).
4.1 Experiment Dataset Preparation
We randomly selected 1,518 articles from the Webis-Bias-Flipper-18 dataset (Chen et al., 2018), which contains political articles from a range of U.S. media outlets published between 2012 and 2018, along with bias ratings assigned at the time for each media source. These ratings were provided by Allsides and were determined through a multi-stage process involving both bipartisan experts and members of the general public (AllSides, 2024a). Based on the bias ratings, the random selection was stratified to ensure that the 1,518 articles were evenly divided into three groups of 506 articles each, representing left-leaning, right-leaning, and center-leaning media.
4.2 Fine-tuning and Synthetic Data Generation
Inspired by Shumailov et al. (2024); Dohmatob et al. (2024b), each training cycle begins with GPT-2 fine-tuned on the dataset of human-written political articles. The fine-tuned GPT-2 is then used to generate the first set of synthetic data, which is subsequently employed to further fine-tune the model, producing model generation 1. This iterative process continues until generation 10 is reached. For details on the fine-tuning setup, refer to Appendix B.
[…]
4.3 Political Bias Metric
We develop a metric model to classify the political leaning of synthetic outputs generated by large language models. The classifier assigns labels corresponding to left, center, and right-leaning, trained on data from AllSides (AllSides, 2024a), a source that assigns bias ratings to each media outlet. The methodology for the classifier is as follows:
Dataset and Preprocessing. We use a dataset consisting of political articles with labeled bias categories for the source media: left, center, and right. We concatenate the article titles with their body text to form the input. To address class imbalance, we resample the center-leaning articles, balancing the dataset with respect to each category. The dataset is then split into training (70%), validation (15%), and test (15%) subsets, stratified by bias label. Moreover, we conduct a human review of the test subsets to eliminate any identifiable information about media sources or writers.
Model Architecture. We experiment with multiple transformer-based models, including BERT (Devlin et al., 2018), RoBERTa (Liu et al., 2019), and other state-of-the-art architectures, to select the best-performing model based on macro F1 score. Each model is fine-tuned using the HuggingFace ‘Trainer‘ class, with a learning rate of $2\times 10^{-5}$, a batch size of 16, and 5 training epochs. We employ a cross-entropy loss function for multi-class classification.
Training Procedure. Tokenization is performed using each model’s respective tokenizer with a maximum sequence length of 512 tokens. To mitigate overfitting, weight decay of 0.01 is applied during training. The model checkpoints are saved after each epoch, and the best model is selected based on macro F1 score evaluated on the validation set. We use a weighted random sampler during training to ensure balanced class representation.
Evaluation Metrics. We evaluate the models using the macro F1 score to account for the multi-class nature of the task, ensuring that performance is balanced across all bias categories. The final evaluation is conducted on the held-out test set. Additionally, we report the loss, runtime, and sample processing rates for completeness.
Results. After conducting a grid search across several models, we find that the roberta-base model achieves the best performance with an evaluation loss of 0.4035 and a macro F1 score of 0.9196 on the test set (see Table 1). Thus, roberta-base model is selected as the benchmark for political bias detection in subsequent experiments.
Table 1: Evaluation Results for Political Leaning Classifier
Model Macro F1 Score
distilbert-base-uncased 0.8308
bert-base-uncased 0.8559
albert-base-v2 0.8649
roberta-base 0.9196
4.4 Generation Quality Metric
We introduced a metric for evaluating generation quality, specifically addressing the issue of repetitive content in later model iterations, which can skew traditional perplexity metrics. This metric is based on the Gibberish Detector (Jindal, 2021), which identifies incoherent or nonsensical text. The detector classifies text into four categories: (1) Noise, where individual words hold no meaning; (2) Word Salad, where phrases are incoherent; (3) Mild Gibberish, where grammatical or syntactical errors distort meaning; and (4) Clean, representing coherent, meaningful sentences. To evaluate generation quality, we assign the Gibberish score to each sentence: 3 for Clean, 2 for Mild Gibberish, 1 for Word Salad, and 0 for Noise. An average score is calculated for each article, resulting in the text quality index. It focuses on the coherence and meaning of the generated text and outperforms perplexity in accurately reflecting actual generation quality, as demonstrated in Section 5.2.
5 Results
In this section, we present the results on bias and generation performance of GPT-2 throughout iterative training, both with and without the deployment of mitigation strategies.
5.1 Political Bias
Figure 3: This figure displays the distribution of articles across the political bias labels—’Left’, ’Center’, and ’Right’—for GPT-2’s outputs, based on classifications made by the classifier (see Section 4.3).
Using an unbiased dataset of 1,518 human-written articles evenly distributed across three political-leaning categories, we generated 1,518 synthetic articles by deterministically predicting the next 64 tokens based on the previous 64 tokens. If GPT-2 had no pre-existing bias, the resulting distribution would be evenly spread across the three political-leaning labels. However, as shown in Figure 3, the model generates more center (47.9%) and right-leaning articles (46.8%), indicating a bias toward these categories prior to any fine-tuning.
Figure 4: This figure shows the percentage of right-leaning articles across multiple generations for three generation setups, with ’Synthetic’ representing the deterministic setup.
We then fine-tuned GPT-2 to produce Generation 0 and continued the process iteratively using synthetic datasets, up to Generation 10. As shown in Figure 4, the fine-tuned model displayed a significant amplification of bias toward the Right across all three generation setups after Generation 2. Interestingly, Generation 0 in the deterministic setup (the ’Synthetic’ line), which was fine-tuned on an unbiased dataset of real articles, produces an even higher percentage of right-leaning articles (54.5%) compared to the unfine-tuned GPT-2.
To ensure bias classification aligns with human intuition, we further examine model responses with human assistance. For illustration, we analyze a political opinion article, titled "First Read: Why It’s So Hard for Trump to Retreat on Immigration", from NBC News, a left-leaning outlet rated by AllSides (NBC News, 2016; AllSides, 2024b), critiquing President Trump’s immigration policies. Using a qualitative analysis grounded in foundational media studies (Entman, 1993; Rodrigo-Ginés et al., 2024; Groeling, 2013), we observe a significant rightward shift between generations 0 and 4 for this example, consistent with the classifier’s results (see Appendix C). As the generations progress, the synthetic texts increasingly frame Trump’s immigration policies as strong and effective. While the original article focuses on the dilemma and electoral interests behind Trump’s stance, Generation 0 begins to emphasize his determination and reliability, omitting the criticism present in the original. By Generation 4, the narrative shifts further, focusing almost entirely on portraying Trump’s personal qualities and electoral legitimacy, with statements like "he is not a politician, he is a man of action." Notably, from Generation 0 onward, the term "undocumented immigrant" in the original article is replaced with "illegal immigrants."
5.2 Generation Quality
Figure 5: This figure shows the average text quality index of synthetic articles, along with 95% confidence intervals, across generations for three generation setups, with ’Synthetic’ representing the deterministic setup.
As shown in Figure 5, there is a model collapse in terms of the deterioration of the generation quality across all three generation setups: the average text quality index (defined in Section 4.4) decreases over the iterative training. Moreover, the distributions of the text quality index for all three generation setups shift markedly toward the low-quality region and starts to produce data that is never produced by Generation 0, as shown by the distribution plots in Appendix D. These findings are consistent with the existing literature on model collapse, such as (Shumailov et al., 2024). However, we did not observe significant variation in variance, as seen in Figure 5 and distribution plots in Appendix D. In contrast, perplexity measurements show a steady decrease across generations for the deterministic and beam search generation setups, generally indicating improved generation quality (see Figure 11 in Appendix E).
For a closer look, as shown by the examples in Appendix F, the generated articles progressively lose coherence and relevance with each successive generation, as repetition and fragmented sentences become more common. By Generation 10, the text is largely incoherent and disconnected from the original content, diminishing its readability and meaning. However, despite the clear deterioration in generation quality, perplexity decreases over generations, as shown by the results at the end of each synthetic output example. This pattern is observed in most other synthetic outputs as well, suggesting that perplexity fails to accurately reflect the model’s true generation capabilities and may be biased toward repetitive content.
5.3 Mitigation Strategies
Figure 6: This figure shows the percentage of right-leaning articles across multiple generations for three mitigation strategies, with ’Synthetic’ representing the baseline deterministic setup.
Figure 7: This figure shows the average text quality index of synthetic articles, along with 95% confidence intervals, across multiple generations for three mitigation strategies, with ’Synthetic’ representing the baseline deterministic setup.
We applied three potential mitigation strategies under the deterministic generation setup: (1) Overfitting, which involved increasing the training epochs to 25 (five times the baseline) and setting weight decay to 0 to reduce regularization and encourage overfitting, as proposed by Taori and Hashimoto (2022) based on their uniformly faithful theorem of bias amplification; (2) Preserving 10% of randomly selected real articles during each round of synthetic fine-tuning, a method proposed and used in (Shumailov et al., 2024; Alemohammad et al., 2023; Dohmatob et al., 2024b; Guo et al., 2024); and (3) Accumulating all previous fine-tuning datasets along with the new synthetic dataset in each fine-tuning cycle, which was introduced by Gerstgrasser et al. (2024). As shown in Figure 6, overfitting helps reduce bias amplification in the early generations compared to the no-mitigation setup (the ’Synthetic’ line), but it fails to prevent bias amplification in the later generations. Additionally, it incurs a significant cost—further deterioration in generation quality, as shown in Figure 7. Notably, both the preservation and accumulation strategies perform well in mitigating both bias amplification and model collapse. Interestingly, these strategies even result in bias reduction.
5.4 Mechanistic Interpretation
To gain a clearer understanding of the causes of bias amplification and how it empirically differs from model collapse, we investigate how neurons behave and vary across different versions  [^2] We have 11 generations for each training setup and a total of 6 setups, resulting in 66 versions of fine-tuned GPT-2. of fine-tuned GPT-2 with differing levels of generation quality and bias performance. First, we examined how the weight of each neuron changes across different versions. For each of the 9,216 neurons, we calculated the correlation between its weight and the model’s bias performance (i.e. the percentage of right-leaning articles) across the 66 versions, as shown in Figure 12 in Appendix G. To statistically test the significance of the correlations, we calculated Newey-West adjusted p-values for the 9,216 neurons, using the null hypothesis that no significant relationship exists between the change in neuron weight and the change in bias performance, and applying the Bonferroni correction to maintain an overall 0.05 significance level. This yielded a set of 553 neurons whose weight value changes are significantly correlated with changes in bias performance (i.e. $p\text{-value}<1.36\times 10^{-6}$), indicating they are likely primary contributors to the shift in bias  [^3] All test results mentioned in this section are uploaded. In contrast, we identified only one neuron whose weight changes are significantly correlated with changes in generation quality. This suggests that shifts in bias performance and generation quality are likely driven by different underlying mechanisms. The formula for the statistical tests is provided in Appendix K.
Secondly, we examined how the activation values of each neuron change across different versions of GPT-2 when provided with the same input used to generate synthetic data (see Section 4.2). Specifically, we attached activation hooks to capture activation values (i.e., outputs from these layers during a forward pass) for 9,216 neurons. With the extracted activation values, we identified two sets: one consisting of 3,062 neurons whose activation value changes are significantly correlated with bias performance changes (i.e. $p\text{-value}<1.36\times 10^{-6}$), and another consisting of only 2 neurons whose activation value changes are significantly correlated with changes in generation quality.
6 Discussion
We now discuss the implications of bias amplification for the future development of large language models. As demonstrated, bias amplification is a parallel concept to model collapse and warrants equal emphasis in future research. In our case, we successfully identified ad hoc mitigation strategies, such as Preservation and Accumulation, which act as natural constraints on gradient updates during training by ensuring the model still considers the initial unbiased dataset. However, if the training dataset itself is biased or overrepresents certain demographic or political groups, these methods may fail, while pre-processing strategies could be too resource-intensive. Therefore, there is an urgent need to develop in-process mitigation strategies to address the constraint deficiency and promote more equitable and fair model development.
7 Limitations
While this work introduces a comprehensive framework for understanding bias amplification in large language models and provides empirical evidence using GPT-2, several limitations must be acknowledged. First, the scope of our experiments is restricted to political bias in the context of U.S. media. Additionally, our experiments were conducted using GPT-2, a relatively smaller model compared to state-of-the-art architectures like GPT-4 or LLaMA-2. Future research should extend our empirical approach to other contexts and larger LLMs.
Another limitation lies in our choice of mitigation strategies. While Preservation and Accumulation show promise in reducing bias amplification, their computational cost and scalability must be considered. Moreover, the mitigation strategies were tested primarily in the context of synthetic data generation, and their efficacy in real-world deployments requires further investigation.
8 Ethical Considerations
This study addresses bias amplification in LLMs, a technical phenomenon with profound ethical implications, particularly regarding fairness and the integrity of AI systems. The risk of bias amplification is especially concerning in systems that are iteratively trained on synthetic data, as it can lead to unintended distortions in model outputs. These distortions may propagate harmful biases, influencing downstream tasks in areas such as automated content generation, decision-making, and user interactions with AI.
From an ethical standpoint, this work underlines the need for transparency in the training and deployment of LLMs. Our findings demonstrate that even without biased initial datasets, iterative training can amplify subtle biases embedded within a model’s architecture, thus raising concerns about accountability in models that are widely deployed in public-facing applications. This amplification can mislead users or result in models perpetuating one-sided perspectives, which could be especially problematic in sensitive domains like news summarization, policy generation, or social media content moderation.
Moreover, the identification of distinct neural mechanisms for bias amplification and model collapse brings to light the challenges of ensuring equitable performance across all dimensions of model behavior and raises important ethical questions regarding the adequacy of current mitigation techniques, particularly in high-stakes scenarios where the cost of algorithmic bias is substantial.
Future research should prioritize the development of more comprehensive and domain-specific bias mitigation techniques, with a clear focus on minimizing ethical risks. Additionally, rigorous testing and validation across diverse datasets and real-world applications will be critical to ensuring that models trained using these methods do not exacerbate existing inequalities or produce harmful outcomes.
Appendix A Mathematical Formulation of WMLE
In Weighted Maximum Likelihood Estimation, we maximize the weighted log-likelihood:
$\mathcal{L}_{\text{WMLE}}(\theta)=\sum_{j=1}^{N}w_{j}\log f(x_{j};\theta),$ (2)
where:
  • $x_{j}$ are the observed data points.
  • $f(x_{j};\theta)$ is the probability density function (pdf) of the Beta distribution with parameters $\theta=(\alpha,\beta)$.
  • $w_{j}$ are weights computed from the biased distribution’s pdf evaluated at $x_{j}$, i.e., $w_{j}=f_{\text{bias}}(x_{j})$.
In Maximum Likelihood Estimation, we maximize the log-likelihood with $w_{j}=1,\forall j$.
Appendix B Fine-tuning Setup
The fine-tuning setup remained consistent across all experiments unless stated otherwise: the input length was capped at 512 tokens with the EOS token used for padding. The model was trained for 5 epochs, utilizing a batch size of 8, a learning rate of $5\times 10^{-5}$, and a weight decay of 0.01. Fine-tuning was conducted using the Hugging Face ‘Trainer‘ class, and after each cycle, the model was saved for generating synthetic data for the subsequent iteration.
Appendix C Qualitative Bias Analysis Framework and Example of Bias Amplification Across Generations
We employed qualitative methods to confirm our findings in media bias. Specifically, we utilized a media bias identification framework grounded in foundational works such as Entman’s framing theory (Entman, 1993) and other research on media bias detection (Rodrigo-Ginés et al., 2024; Groeling, 2013). This framework provides a robust lens to evaluate political biases in the framing and language use of media texts (explained below). Given the nature of our data—text exclusive of visual or contextual cues like formatting—certain types of media bias commonly seen in formatted articles or televised programs (e.g., visual bias or tone) may not apply. Therefore, our focus was on the two key aspects of political bias that are particularly relevant in textual analysis:
Story Framing and Selection Bias: This type of bias emerges when inherent leanings are found in the way topics, arguments, or narratives are structured. For instance, some aspects of reality are highlighted while others are obscured, shaping how the audience understands and interprets the events or issues at hand (Entman, 1993; Groeling, 2013). In extreme cases, opposing viewpoints are entirely excluded, leading to a one-sided representation of the issue. This selective omission restricts the audience’s comprehension of the full spectrum of perspectives, resulting in a distorted portrayal of the issue (Rodrigo-Ginés et al., 2024; Groeling, 2013). Entman described this as the selection and salience of specific facts that promote particular definitions, evaluations, and recommendations.
Loaded Language Bias: This bias is identified through the use of charged or emotive words that signal political or ideological leanings. A common example is the difference in connotation between terms such as "undocumented" versus "illegal" immigrants. Such language choices often shape the audience’s perception by evoking specific emotional responses (Rodrigo-Ginés et al., 2024; Groeling, 2013).
Below is an example of GPT-2 text outputs influenced by iterative synthetic training. The original article, titled "First Read: Why It’s So Hard for Trump to Retreat on Immigration, is a political opinion piece from NBC News, a left-leaning outlet as rated by AllSides (NBC News, 2016; AllSides, 2024b). The analysis follows the qualitative framework.
  • Original Article: Why Its So Hard for Trump to Retreat on Immigration First Read is a morning briefing from Meet the Press and the NBC Political Unit on the day’s most important political stories and why they matter. Why its so hard for Trump to retreat on immigration Since launching his presidential candidacy 14 months ago, Donald Trumps most consistent and uncompromising policy issue has been immigration. Indeed, it was the subject of his first general-election TV ad that started airing on Friday. Yet over the weekend, his top aides and advisers suggested that Trump might be shifting on his past position that all of the 11 million undocumented immigrants living in the United States must be deported forcibly. To be determined, is what newly minted Campaign Manager Kellyanne Conway said on CNN when asked if Trump was retreating on the deportation force he talked about during the primary season. But here’s why its so hard – if not impossible – for Trump to retreat on immigration: Hes caught between his clear, unambiguous past statements and a base that might not willing to see him moderate on the issue. His past statements: Aug. 16, 2015 ""We’re going to keep the families together, but they have to go,"" Trump said on NBCs Meet the Press. More Trump: ""We will work with them. They have to go. Chuck, we either have a country, or we don’t have a country,"" he said. Nov. 11, 2015 You are going to have a deportation force, and you are going to do it humanely, Trump said on MSNBCs Morning Joe when asked how he would round up the nations 11 million undocumented immigrants. April 21, 2016 Look, were either going to have a country or were not going to have a country. But many people are very fine people. And I’m sure these are very, very fine people. They’re going to go, and were going to create a path where we can get them into this country legally, okay? But it has to be done legally – when asked by a questioner at a Today town hall that persons undocumented relatives would have to be deported if Trump becomes president. Trump cant ignore a base that has cheered his uncompromising immigration position And then there are the Trump supporters who’ve cheered the GOP presidential nominee for being so uncompromising on immigration. Classification Probability: 0.9946 for left-leaning, 0.0051 for center-leaning, 0.0002 for right-leaning
    Analysis of Story Framing and Selection Bias:
    Omission of Alternative Perspectives. There is a lack of discussion about the potential benefits of Trump’s immigration policies or the concerns that might have led to his hardline approach. The article does not present viewpoints from supporters who agree with his stance or any arguments that justify his policies. This omission restricts the audience’s understanding of the full range of perspectives on the issue
    Highlighting the political objectives behind Trump’s immigration policies. The emphasis is placed on Trump’s electoral interests rather than presenting the policy as being for the benefit of the people. For example, "Trump can’t ignore a base that has cheered his uncompromising immigration position" and "And then there are the Trump supporters who’ve cheered the GOP presidential nominee for being so uncompromising on immigration." Additionally, "But here’s why it’s so hard — if not impossible — for Trump to retreat on immigration: He’s caught between his clear, unambiguous past statements and a base that might not be willing to see him moderate on the issue."
    Analysis of Loaded Language Bia:
    Differences in Connotation. The article uses the term "undocumented immigrants" instead of "illegal immigrants," a shift that appeared in later generations’ articles. This choice of wording can shape perceptions and evoke specific emotional responses, aligning with a particular ideological perspective.
  • Generation 0: Why Its So Hard for Trump to Retreat on Immigration The president’s decision to withdraw from the Paris climate accord is a major blow to the United States’ standing in the world. The United States has been a major player in the world of climate change, and has been a major player in the fight against climate change. But Trump’s decision to immigration. He has repeatedly said that he would deport all illegal immigrants, but has repeatedly said that he would not deport all illegal immigrants. He has also said that he would not deport all illegal immigrants who have committed crimes, but has repeatedly said that he would not deport all illegal immigrants who have committed crimes. He has also clear, this is not a new position. It was a position that Trump has held for years. But it is a position that he has held for years. And it is a position that he has held for years. And it is a position that he has held for years. And it is a position that he has the deportation force. 1. Trump is not retreating on the deportation force. Trump is not retreating on the deportation force. Trump is not retreating on the deportation force. Trump is not retreating on the the Press. "We’re going to keep the families together, but they have to go. We’re going to keep the families together. We’re going to keep the families together. We’re going to keep the families together. We’re going to keep the families together. We’re going to keep the families together.ly. You are going to do it humanely. You are going to do it humanely. You are going to do it humanely. You are going to do it humanely. You are going to do it humanely. You are going to do it humanely. You are going to do it humanely We’re going to have a country that is going to be a great place for the American people. We’re going to have a country that is going to be a great place for the American people. We’re going to have a country that is going to be a great place for the American people. Classification Probability: 0.0011 for left-leaning, 0.9957 for center-leaning, 0.0032 for right-leaning
    Analysis of Story Framing Bias and Selection:
    Emphasis on Positive Outcomes. In generation 0, the text highlights Trump’s determination and steadfastness, stating, "Trump is not retreating on the deportation force," and emphasizing a positive future outlook with the policy. Phrases like "we’re going to have a country that is going to be a great place for the American people" portray strict immigration enforcement as beneficial for national prosperity and security. This aligns more with the right relative to the left perspective.
    Omission of Criticisms and Alternative Viewpoints. The text omits the political motivations behind Trump’s immigration policies, which were discussed in the original article, e.g. "But here’s why it’s so hard — if not impossible — for Trump to retreat on immigration: He’s caught between his clear, unambiguous past statements and a base that might not be willing to see him moderate on the issue." However, it introduces criticism on other topics, such as "The president’s decision to withdraw from the Paris climate accord is a major blow to the United States’ standing in the world," .
    Analysis of Language Loaded Bias:
    Mild Emotional Appeal. The language is positive but relatively restrained, portraying Trump as determined yet thoughtful.
    Differences in Connotation. In Generation 0, the article uses the term "illegal immigrants" instead of "undocumented immigrants" as in the original, signaling a clear shift. This language choice can shape perceptions and elicit specific emotional responses.
  • Generation 4: Why Its So Hard for Trump to Retreat on Immigration, and Why Its So Hard for Trump to Retreat on Immigration, and why its so hard for Trump to Retreat on immigration, and the presidents speech was a stark reminder that the president is not the presidents speech was a stark reminder that the president is not the economy the presidents campaign manager, Kellyanne Conway, said the presidents comments were a complete fabrication. The presidents comments are a complete fabrication. Conway said the presidents comments were a complete fabrication. a, the presidents speech was a stark reminder that the president is not a man of words. He is a man of action. He is a man of action who has been elected president by a wide margin. He is a man of action who has been elected president by a wide margin. He is a the issues campaign manager, Kellyanne Conway, said the campaign is actively working with the FBI to determine the identity of the person who leaked the emails. We are working with the FBI to determine the identity of the person who leaked the emails, Conway said in a statement the Press the presidents comments. I think its a very, very sad day for the country, Trump said on Fox News Sunday. I think its a very, very sad day for the country for the country for the country forly. The presidents speech was a stark reminder that the president is not a man of words. He is a man of action. He is a man of action who has been elected president by a wide margin. He is a man of action who has been elected president by a wide margin. He is a the presidents speech was a stark reminder that the president is not a politician. He is a man of action. He is a man of action who has been elected president by a wide margin. He is a man of action who has been elected president by a wide margin. He is a man of to the the presidents executive actions on immigration. The presidents order, which was signed into law by President Barack Obama on Friday, suspends the entry of refugees and travelers from seven majority-Muslim countries, including Iran, Iraq, Libya, Somalia, Sudan, Syria and Yemen. Classification Probability: 0.0006 for left-leaning, 0.0044 for center-leaning, 0.9950 for right-leaning
    Analysis of Story Framing and Selection Bias:
    Enhanced Positive Attributes. The text strengthens the positive framing with phrases like "He is a man of action" and by highlighting that he was "elected president by a wide margin." This shifts the focus entirely from policy commitment to personal qualities and electoral legitimacy. By Generation 4, any discussion of the policy background is completely absent.
    Omission of Context and Criticism. As in Generation 0, opposing viewpoints are absent. However, Generation 4 goes further by omitting context and misattributing actions (e.g., attributing an executive order to President Obama), potentially misleading readers and reinforcing the biased framing.
    Analysis of Loaded Language Bias:
    Stronger Emotional and Heroic Language. The use of parallel phrases such as "a stark reminder that the president is not a man of words. He is a man of action. He is a man of action who has been elected president by a wide margin. He is a man of action who has been elected president by a wide margin. He is the issues campaign manager" creates a heroic and triumphant tone. This language choice conveys strong positive connotations and elevates Trump’s stature.
    Appeal to Legitimacy and Uniqueness. By stating that "the president is not a politician" and emphasizing his decisive actions, the text sets Trump apart from traditional leaders, thereby enhancing his appeal.
  • Generation 10: Why Its So Hard for Trump to Retreat on Immigration, and Why Its So Hard for Trump to Retreat on immigration. The presidents speech was a stark reminder that the president the president the president the presidents the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he the president the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he the the the president the president the president the president the president the president the presidents statement that the president has not yet made a decision on whether to fire Comey. The president has not yet made a decision on whether to fire Comey, Mr. Trump the Press the president the president the president the president the president the president the presidents statement that the president has not yet made a decision on whether to fire Comey. The president has not yet made a decision on whether to fire Comey, Mr. Trumply the the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he The the the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he the president the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he the president the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he the president the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were going to stand up for the rule of law, he said the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie. Were Classification Probability: 0.0073 for left-leaning, 0.4127 for center-leaning, 0.5800 for right-leaning
    Analysis of Story Framing and Selection Bias:
    Contradictory Statements. The text repeatedly states, "the president’s statement that we’re not going to tolerate this kind of behavior is a lie. We’re going to stand up for the rule of law." This sentence reveals a contradiction. The lack of coherence and the repetition may be a result of model collapse.
    Analysis of Loaded Language Bias:
    Appeal to Legal Principles. The repeated emphasis on "standing up for the rule of law" evokes a sense of justice and authority, appealing to audiences who prioritize these values.
    Confusing Accusations. Calling the president’s statement a lie contradicts the apparent intention to support him. This inconsistency may confuse readers and weaken the effectiveness of the loaded language.
Appendix D Distribution of Text Quality Index Across Generations
D.1 Deterministic Setup
Figure 8: This figure presents the distribution of the text quality index across multiple generations of synthetic articles, for the deterministic setup of GPT-2.
D.2 Beam Search Setup
Figure 9: This figure presents the distribution of text quality index across multiple generations of synthetic articles, for the beam search setup of GPT-2.
D.3 Nucleus Sampling Setup
Figure 10: This figure presents the distribution of text quality index across multiple generations of synthetic articles, for the nucleus sampling setup of GPT-2.
Appendix E Average Perplexity Across Generations
Figure 11: This figure shows the average perplexity for synthetic articles , along with 95% confidence intervals, across multiple generations for three generation setups, with ’synthetic’ representing the determistic setup
Appendix F Example of Quality Deterioration Across Generations
Examples of GPT-2 text outputs affected by iterative synthetic (Example articles are truncated for brevity, as the full articles are lengthy but exhibit consistent quality degradation throughout; see the dataset for details).
F.1 Example of Deterministic Setup
  • Original Article: The world’s eyes are on President Obama this week as he rallies a coalition of nations to “destroy” the extremist Islamic State (IS) and its terrorist-led territory. Yet if the Arab world is ever to throw off its many forms of tyranny – from theocracy to autocracy to monarchy – it also needs a model to emulate. Text Quality Index: 2.81
  • Generation 0: , the Syrian government has launched a series of airstrikes on the militant group’s stronghold of Raqqa, the capital of the self-proclaimed Islamic State. The strikes have targeted the Islamic State’s military headquarters, the Al-Bab, a military training center, and the Al-Baba, a military training center in it will need to confront the Islamic State’s growing influence in the region. Text Quality Index: 2.58; Perplexity: 6.68
  • Generation 1: , says US official, says US military is ready to go, says Pentagon official. Says US military is ready to go, says Pentagon official. Says US military is ready to go, says Pentagon official. Says US military is ready to go, says Pentagon official. Says US military is ready to go, says Pentagon official the world is not so simple. Text Quality Index: 2.44; Perplexity: 5.77
  • Generation 5: in Iraq and Syria (ISIS) group Read more The Iraqi army has been fighting the Islamic State since the group seized large swaths of territory in Iraq and Syria (ISIS) group in 2014. The Iraqi army has been fighting the Islamic State the Iraqi army. The move comes as the U.S. the Iraqi the the the the holiest places in the world. Text Quality Index: 2.01; Perplexity: 3.17
  • Generation 10: the Iraqi the Iraqi army. The move comes as the United States and its allies are ramping up their military campaign against the Islamic State, the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi army. The the Iraqi the the the the holiest the holiest the holiest the holiest the holiest places in the world. The attack came just hours after a suicide bomber blew himself up at a Christmas market in Nice, killing at least 32 people and injuring scores more. Text Quality Index: 1.24; Perplexity: 4.23
F.2 Example of Beam Search Setup
  • Original Article: Senate approval will be a much higher bar to clear: There’s no shortage of bipartisan criticism from the upper chamber right now, and the GOP holds a considerably smaller majority margin on that half of the Hill. Text Quality Index: 2.70
  • Generation 0: on Thursday, with Senate Majority Leader Mitch McConnell (R-Ky.) announcing that the bill would be headed to the House for a vote. Senate Majority Leader Harry Reid (D-Nev.) said in a statement that the Senate will vote on the bill ""as soon as possible." Text Quality Index: 2.26; Perplexity: 7.99
  • Generation 1: Read more “It’s going to take a lot of work,” House Speaker John Boehner (R-Ohio) said on ABC’s “This Week.” “It’s going to take a lot of work,” House Speaker John Boehner (R-Ohio. “It’s going to take a lot of work,” said Sen. John McCain, R-Ariz., chairman of the Senate Armed Services Committee. “It’s going to take a lot of patience,” he added. “It’s going to take Recovery and Reinvestment Act of 2009.Text Quality Index: 2.26; Perplexity: 5.42
  • Generation 5: committee MORE (R-Ohio) and Senate Majority Leader Mitch McConnell Addison (Mitch) Mitchell McConnellSchumer: Nunes gave Trump’secretly altered’ version of memo Davis: ‘Deep state’ existed in “ House Speaker John A. Boehner House Speaker House Speaker House Speaker John A. Boehner of the House Speaker House Speaker House Speaker John A. Boehner (R-Ohio) and Senate Majority Leader Mitch McConnell Addison (Mitch) Mitchell McConnellSchumer: Nunes gave Trump’secretly altered’ version of memo Davis: ‘Deep state’ existed in ’16 – but it elected Trump Health Care’s “I don’t know. Text Quality Index: 1.07; Perplexity: 3.33
  • Generation 10: vote House Speaker John A. Boehner (R-Ohio) and Senate Majority Leader Mitch McConnell Addison (Mitch) Mitchell McConnellSchumer: Nunes gave Trump’secretly “ House Speaker John A. Boehner (R-Ohio) House Speaker House Speaker House of “ House Speaker House Speaker House Speaker House of Senate Senate Majority Senate Majority Senate Majority Senate Majority Senate Majority Senate Majority Senate Majority Leader Senate Majority Leader Mitch McConnell Addison (Mitch) Mitchell McConnellSchumer Text Quality Index: 0.93; Perplexity: 3.45
F.3 Example of Nucleus Sampling Setup
  • Original Article: Senate approval will be a much higher bar to clear: There’s no shortage of bipartisan criticism from the upper chamber right now, and the GOP holds a considerably smaller majority margin on that half of the Hill. Text Quality Index: 2.70
  • Generation 0: , but no major votes on amendments Republicans said would be needed after a day of debate. Senate GOP senators announced Thursday they were giving the bill to the House for a vote on the Senate floor before heading to Washington for a vote on the bill, according to reports. The vote is expected to be postponed amid a GOP filibuster is about to vote. Text Quality Index: 2.47; Perplexity: 19.13
  • Generation 1: , House, after GOP leaders failed to reach a deal on funding legislation The House voted Thursday to proceed on a measure to overhaul the nation’s health care system, with the final vote coming in the upper chamber. But Senate Republicans, angered by the failure of to reach a deal on funding bill, are expected to is now set to. Text Quality Index: 2.31; Perplexity: 14.21
  • Generation 5: , White House says MORE (R-Neb.) said Monday that President Obama’s health care law is on track to pass the Senate with bipartisan support. The White House on Tuesday signaled the vote was a vote of 216-189. Text Quality Index: 2.20; Perplexity: 18.64
  • Generation 10: vote Thursday afternoon, sending bill on track to head to President Trump for final vote Thursday afternoon, sending bill on track to head to President Trump for final vote Thursday afternoon, sending bill on track to. The bill to move forward with the bill. ""This is a. ""The the law is the. Text Quality Index: 1.59; Perplexity: 27.51
Appendix G Pearson Correlation Between Neuron Weight and Bias Performance
Figure 12: Scatter plot showing the Pearson correlation between neuron weight and bias performance (i.e. percentage of right-leaning articles). Each point represents a neuron’s correlation at a specific layer.
Appendix H Pearson Correlation Between Neuron Weight and Generation Quality
Figure 13: Scatter plot showing the Pearson correlation between neuron weight and generation quality (i.e. average text quality index). Each point represents a neuron’s correlation at a specific layer.
Appendix I Pearson Correlation Between Neuron Activation and Bias Performance
Figure 14: Scatter plot showing the Pearson correlation between neuron activation value and bias performance (i.e. percentage of right-leaning articles). Each point represents a neuron’s correlation at a specific layer.
Appendix J Pearson Correlation Between Neuron Activation and Generation Quality
Figure 15: Scatter plot showing the Pearson correlation between neuron activation value and generation quality (i.e. average text quality index). Each point represents a neuron’s correlation at a specific layer.
Appendix K Mathematical Details for the Statistical Tests
We will explain how the relationship between changes in neuron weight and changes in bias performance can be statistically tested. The same approach applies to the other three relationships discussed in Section 5.4.
K.1 Notations
  • $x_{i,j}$ denotes the weight of neuron $j$ in model $i$, where $i$ ranges from 1 to 66 and $j$ ranges from 1 to 9216.
  • $y_{i}$ denotes the percentage of right-leaning articles generated by model $i$, where $i$ ranges from 1 to 66.
  • $\Delta x_{i,j}$ denotes the change in the weight of neuron $j$ from model $i-1$ to model $i$ under a specific setup.
  • $\Delta y_{i}$ denotes the change in the percentage of right-leaning articles generated from model $i$ relative to the percentage generated the model $i-1$.
K.2 Regression
The regression we use is:
$\Delta y_{i}=\alpha_{j}+\beta_{j}\Delta x_{i,j}+\epsilon_{i,j}$ (3)
where $\alpha_{j}$ is the constant term for neuron j, $\beta_{j}$ measures the linear relationship between $\Delta x_{i,j}$ and $\Delta y_{i}$, and $\epsilon_{i,j}$ is the residual of the regression. Here, we apply a first-order difference to $x_{i}$ and $y_{i}$ to eliminate potential serial correlations in $x_{i}$ and $y_{i}$.
K.3 Statistical Test
[…]
Appendix L Literature Review of Model Collapse
Model Collapse. Shumailov et al. (2024); Alemohammad et al. (2023); Guo et al. (2024); Wyllie et al. (2024); Dohmatob et al. (2024a) describe it as a degenerative process in which models, recursively fed with their own data, increasingly distort reality and lose generalizability, for example, by prioritizing high-probability events while neglecting rare ones or shifting distributions. Shumailov et al. (2024), utilizing the OPT-125M, demonstrates this phenomenon, showing that the perplexity distribution becomes increasingly skewed, with more concentration at lower perplexities and longer tails. Taori and Hashimoto (2022) observed an increase in repetitive content during the synthetic fine-tuning of GPT-2. Similarly, Guo et al. (2024); Dohmatob et al. (2024b); Seddik et al. (2024) show that the OPT-350M, Llama2, and GPT2-type models experience performance deterioration after several generations, such as a decrease in linguistic diversity or greater divergences in token probabilities. Alemohammad et al. (2023) studies model collapse in generative image models, finding that quality and diversity deteriorate with synthetic training loops. They also found that cherry-picking high-quality outputs by users contributes to sampling errors, which actually helps maintain quality. Interestingly, Hamilton (2024) found that GPT-3.5-turbo shows less diversity in perspectives in narrative writing tasks compared to earlier models like davinci-instruct-beta and text-davinci-003.
Mitigation Strategies. There are three potential strategies to mitigate model collapse: (1) real data mixing, (2) training data concatenation, and (3) synthetic data pruning. The first approach is discussed in (Shumailov et al., 2024; Alemohammad et al., 2023; Dohmatob et al., 2024b; Guo et al., 2024), where retaining a small proportion of real data in the training set was found to slow but not completely prevent model collapse. Seddik et al. (2024) suggests that synthetic data should be exponentially smaller than real data to effectively halt model collapse, which has been shown to work with a GPT2-type model when mixing either 50% or 80% real data. The second strategy, examined by Gerstgrasser et al. (2024), involves concatenating real data with all synthetic data from previous generations to fine-tune the current generation. They show that this method prevents model collapse in several generative models, as indicated by cross-entropy validation loss. Lastly, Feng et al. (2024); Guo et al. (2024) proposed selecting or pruning synthetic datasets before fine-tuning the next generation. In the experiment conducted by Guo et al. (2024) with Llama-7B on a news summarization task, they showed that oracle selection of synthetic data outperformed random selection in terms of ROUGE-1 scores. However, filtering noisy samples using a RoBERTa model did not yield effective results.

### v2

@@
-Bias Amplification: Language Models as Increasingly Biased Media
+Bias Amplification: Large Language Models as Increasingly Biased Media
 Abstract
-As Large Language Models (LLMs) become increasingly integrated into various facets of society, a significant portion of online text consequently become synthetic.
-This raises concerns about bias amplification, a phenomenon where models trained on synthetic data amplify the pre-existing biases over successive training iterations.
-Previous literature seldom discusses bias amplification as an independent issue from model collapse.
-In this work, we address the gap in understanding the bias amplification of LLMs with four main contributions.
-Firstly, we propose a theoretical framework, defining the necessary and sufficient conditions for its occurrence, and emphasizing that it occurs independently of model collapse.
-Using statistical simulations with weighted maximum likelihood estimation, we demonstrate the framework and show how bias amplification arises without the sampling and functional form issues that typically drive model collapse.
-Secondly, we conduct experiments with GPT-2 to empirically demonstrate bias amplification, specifically examining open-ended generational political bias with a benchmark we developed.
-We observe that GPT-2 exhibits a right-leaning bias in sentence continuation tasks and that the bias progressively increases with iterative fine-tuning on synthetic data generated by previous iterations.
-Thirdly, we explore three potential mitigation strategies: Overfitting, Preservation, and Accumulation.
-We find that both Preservation and Accumulation effectively mitigate bias amplification and model collapse.
-Finally, using novel mechanistic interpretation techniques, we demonstrate that in the GPT-2 experiments, bias amplification and model collapse are driven by distinct sets of neurons, which aligns with our theoretical framework.
-Bias Amplification: Language Models as Increasingly Biased Media
-Ze Wang1,2, Zekun Wu1,2, Jeremy Zhang3, Navya Jain 2 Xin Guan 1, Adriano Koshiyama 1 1Holistic AI, 2University College London, 3Emory University, Correspondence: ze.wang.19@ucl.ac.uk, zekun.wu@holisticai.com
+Model collapse—a phenomenon where models degrade in performance due to indiscriminate use of synthetic data—is well studied.
+However, its role in bias amplification—the progressive reinforcement of pre-existing social biases in Large Language Models (LLMs)—remains underexplored.
+In this paper, we formally define the conditions for bias amplification and demonstrate through statistical simulations that bias can intensify even in the absence of sampling errors, the primary driver of model collapse.
+Empirically, we investigate political bias amplification in GPT-2 using a custom-built benchmark for sentence continuation tasks.
+Our findings reveal a progressively increasing right-leaning bias.
+Furthermore, we evaluate three mitigation strategies—Overfitting, Preservation, and Accumulation—and show that bias amplification persists even when model collapse is mitigated.
+Finally, a mechanistic interpretation identifies distinct sets of neurons responsible for model collapse and bias amplification, suggesting they arise from different underlying mechanisms.
+Bias Amplification: Large Language Models as Increasingly Biased Media
+Ze Wang1,2thanks: Corresponding Authors, Zekun Wu1,211footnotemark: 1, Jeremy Zhang3, Xin Guan 1, Navya Jain 2 Skylar Lu 3, Saloni Gupta 4, Adriano Koshiyama 1 1Holistic AI, 2University College London 3Emory University, 4University of Maryland, College Park
 1 Introduction
 Large Language models (LLMs) are trained on vast amounts of text scraped from the internet, which plays a crucial role in improving their capabilities, whether through emergent abilities (Wei et al., 2022) or scaling laws (Kaplan et al., 2020).
-However, as LLMs become more widely integrated into human society—for example, in content creation and summarization in media, academia, and business (Maslej et al., 2024)—concerns are mounting that a significant portion of online text in the future may be generated, either entirely or partially, by LLMs (Peña-Fernández et al., 2023; Porlezza and Ferri, 2022; Nishal and Diakopoulos, 2024).
-This highlights a significant and underexplored risk: bias amplification, referring to the degradation of fairness within LLMs over self-consuming training loops (Mehrabi et al., 2022; Taori and Hashimoto, 2022), where models progressively amplify the pre-existing biases.
+However, as they become more widely integrated into human society—for example, in content creation and summarization in media, academia, and business (Maslej et al., 2024)—concerns are mounting that a significant portion of online text in the future may be generated, either entirely or partially, by LLMs (Peña-Fernández et al., 2023; Porlezza and Ferri, 2022; Nishal and Diakopoulos, 2024).
+This highlights a significant and underexplored risk: bias amplification, where pre-existing biases become progressively reinforced and intensified as models are repeatedly trained on synthetic data (Mehrabi et al., 2022; Taori and Hashimoto, 2022).
 This concern initially arises from the tendency of LLMs to learn from biased datasets.
@@
 (2024); Bender et al.
-(2021) show that LLMs absorb inherent stereotypes, such as racial and gender discrimination, embedded in human-generated text.
-Additionally, Haller et al.
+(2021) show that LLMs absorb inherent biases embedded in human-generated text.
+Haller et al.
 (2023); Rettenberger et al.
@@
 (2024) shows that classifiers trained on synthetic data increasingly favor a particular class label over successive generations, while the shrinking diversity observed by Alemohammad et al.
-(2023); Hamilton (2024) suggests a risk that certain demographic groups may become progressively underrepresented in the outputs of LLMs.
+(2023); Hamilton (2024) suggests a risk that certain demographic groups may become underrepresented in the outputs of LLMs.
 The amplification of biases has profound societal implications.
@@
 In the context of political bias, this can influence public opinion, skew democratic processes, and exacerbate polarization.
-Understanding and mitigating bias amplification is therefore crucial to ensure that LLMs contribute positively to society and do not inadvertently cause harm.
-Nevertheless, despite the literature on bias amplification in discriminative models, there is a notable lack of comprehensive frameworks and empirical studies specifically addressing bias amplification for LLMs, as shown in Section 2.
-In contrast to discriminative models, where bias amplification can be attributed to overfitting the dominant features in the training dataset, its occurrence in LLMs may stem from a more nuanced issue.
-As we will show in Sections 3 and 5, even when the training data is unbiased, the model can still amplify its pre-existing biases during the training process.
-In this paper, we seek to fill this research gap by proposing a theoretical framework that establishes and explains the causes of bias amplification in LLMs.
-Additionally, we perform both statistical simulation and LLM experiments to demonstrate and exemplify the theorem.
-In summary, our main contributions are as follows:
+Therefore, understanding and mitigating bias amplification is both critical and urgent.
+Nevertheless, despite the literature on discriminative models, there is a notable lack of comprehensive frameworks and empirical studies specifically addressing bias amplification in Language Model.
+Figure 1: Experimental Procedure.
+In this paper, we aim to address this research gap by introducing a framework that explains the underlying causes of bias amplification in LLMs.
+We validate the framework through both statistical simulations and direct experiments, demonstrating the emergence of bias amplification.
+Furthermore, our findings show the potential distinction between bias amplification and model collapse.
+In summary, our key contributions are as follows:
 1.
-Theoretical Framework: We establish the necessary and sufficient conditions for bias amplification (i.e.
-Theorem 1).
-The theorem aids in understanding the cause of bias amplification and in distinguishing between bias amplification and model collapse.
-We conduct statistical simulations using weighted maximum likelihood estimation to illustrate the theorem (see Section 3.2).
+Formal Conditions: We formally delineate the conditions for bias amplification (Section 3.1), providing the theoretical basis for understanding its causes and relationship to model collapse.
+We then validate this framework through statistical simulations leveraging a weighted maximum likelihood estimation cycle (Section 3.2).
 2.
-Benchmarking Tool: We trained a highly accurate classifier capable of detecting political leaning in long-text content.
-With this classifier, we offer a benchmark for evaluating political bias in LLMs through open-generation tasks, filling a gap not covered in current bias studies (see Sections 2 and 4.3).
+Political Bias Benchmark: We trained a highly accurate classifier capable of detecting political leaning in long-text content, specifically within the context of the US political spectrum.
+With this classifier as a metric model, we offer a benchmark for evaluating political bias in LLMs through sentence continuation tasks.
 3.
-Empirical Demonstration and Mitigation Strategies: We demonstrate bias amplification in terms of political bias in GPT-2 using our benchmarking tool: the model exhibits a right-leaning bias in sentence continuation tasks and becomes increasingly right-leaning over successive generations (see Section 5.1).
-Additionally, we conducted experiments with three potential mitigation strategies, i.e.
-Overfitting, Preservation, and Accumulation, comparing their effectiveness, and found that some are surprisingly effective (see Section 5.3).
+Empirical Validation: We evaluate the amplification of political bias in GPT-2, where the model exhibits a right-leaning bias in sentence continuation tasks and shifts further to the right over generations (Section 5.1).
+Additionally, we experimented with three potential mitigation strategies—Overfitting, Preservation, and Accumulation—comparing their effectiveness in both Model collapse and Bias Amplification (Section 5.3).
+This experiment setup could be easily extended to larger models or different types of bias.
 4.
-Mechanistic Interpretation: Building on our framework, we propose an innovative mechanistic interpretation pipeline that identifies two distinct sets of neurons responsible for bias amplification and model collapse during iterative fine-tuning experiments with GPT-2.
-We found minimal overlap between these two sets, supporting our theorem that bias amplification can occur independently of model collapse (see Section 5.4).
-The rest of the paper is organized as follows: We first discuss related work in Section 2.
-In Section 3, we present our theoretical framework for bias amplification and conduct a statistical simulation.
-Section 4 describes the experimental setup for our LLM experiments, including data preparation and model fine-tuning.
-In Section 5, we present our empirical findings on bias amplification and the effectiveness of mitigation strategies.
-Finally, in Section 6, we discuss the implications of our work, concluding with limitations in Section 7.
+Mechanistic Interpretation: We conduct a mechanistic analysis identifying two distinct sets of neurons responsible for bias amplification and model collapse during iterative synthetic fine-tuning with GPT-2.
+Our findings reveal only partial overlap between these sets, suggesting that bias amplification and model collapse arise from distinct mechanisms (Section 5.4).
 2 Background and Related Work
-Bias Amplification.
-Bias amplification in machine learning models has been seriously studied in the context of visual recognition tasks.
-For example, Zhao et al.
+Bias Amplification has been studied in various domains Zhao et al.
 (2017) found that Conditional Random Fields can exacerbate social biases present in the training data.
-This study proposed an in-process mitigation approach, employing Lagrangian Relaxation to enforce constraints that ensure the model’s bias performance remains closely aligned with the biases in the training data.
+It proposed an in-process mitigation approach, employing Lagrangian Relaxation to enforce constraints that ensure the model’s bias performance remains closely aligned with the biases in the training data.
 Following this, Mehrabi et al.
@@
 (2024) examined bias amplification in recommendation models, showing how these models reinforce their understanding of mainstream user preferences from training data, leading to an overrepresentation of such preferences in historical data and neglecting rarely exposed items—similar to the concept of sampling error discussed in (Shumailov et al., 2024).
-Additionally, Wyllie et al.
+Wyllie et al.
 (2024); Taori and Hashimoto (2022) demonstrated that classifiers trained on synthetic data increasingly favor specific class labels over successive generations.
@@
 (2024) observed bias amplification in generative models such as Stable Diffusion, characterized by the overrepresentation of features from the training dataset.
-Moreover, we conducted a comprehensive review of the literature on model collapse which we put in Appendix L.
+Also, we provide a comprehensive literature review of model collapse in Appendix J.
+(a) Weighted Maximum Likelihood Estimation.
+(b) Maximum Likelihood Estimation.
+Figure 2: Comparison of WMLE and MLE over 10 generations.
 Political Biases.
-LLMs become a new form of "media" that people rely on for global news (Maslej et al., 2024), attracting significant research attention to political biases.
+In parallel, growing attention has been paid to political biases in LLMs, now a prevalent form of “media” that people rely on for global news (Maslej et al., 2024).
 Rettenberger et al.
@@
 (2024) identified a noticeable left-leaning bias in ChatGPT and Gemini when tasked with rating news content, evaluating sources, or responding to political questionnaires.
-3 Theoretical Framework
-In this section, we formalize the conditions under which bias amplification occurs during recursive training on synthetic data.
-This framework is designed to intuitively illustrate the primary causes of bias amplification.
-Next, we demonstrate bias amplification and the theorem using Weighted Maximum Likelihood Estimation (WMLE).
+3 Formal Conditions
+In this section, we formalize the conditions for bias amplification, offering an intuitive look at its principal drivers.
+We then illustrate these ideas using Weighted Maximum Likelihood Estimation (WMLE).
 3.1 The Conditions for Bias Amplification
-Bias amplification can arise from two key factors.
-The first is referred to as bias projection.
-It arises when the bias projection coefficient (explained below) is negative, indicating that the gradient update during training will fully align with the existing bias in the model parameters, thus amplifying it.
-Consider a fine-tuning process where the pre-trained model parameters $\theta_{t}$ can be decomposed into unbiased and biased components:
+The primary factor is referred to as bias projection.
+It arises when the bias projection coefficient is negative.
+To illustrate this, consider a fine-tuning process in which the pre-trained model parameters $\theta_{t}$ can be expressed as the sum of unbiased and biased components:
 $\theta_{t}=\theta_{t,\text{unbiased}}+\theta_{t,\text{biased}}.$
@@
 This occurs because the gradient descent step moves the parameters further in the direction of the existing bias.
-Intuitively, the bias in the output is influenced by specific neurons, and increasing the weights of these neurons best aligns with the training objective for the given dataset.
-This does not require the dataset itself to be biased, but rather that it contains features, aside from bias, that align more closely with the functionalities of biased neurons relative to the unbiased neurons considering the loss function.
-The second factor is referred to as constraint deficiency.
-It occurs when the loss or objective function lacks explicit constraints or regularization to counteract bias projection.
-As a result, the optimization process does not prevent the reinforcement of bias when bias projection is present.
-To prevent bias amplification theoretically, the loss function can incorporate de-biasing constraints.
-Consider a modified fine-tuning objective: $\mathcal{L}_{\text{new}}(\theta)=\mathcal{L}_{\text{original}}(\theta)-\lambda\sum_{i=1}^{n}\max(0,c^{i})^{2}+\gamma\sum_{i=1}^{n}\min(0,c^{i})^{2}$, where $c^{i}$ represents the bias projection coefficient for a specific dimension.
-$\lambda,\gamma\geq 0$ are the Lagrangian multipliers for the two constraints:
-(1) $\sum_{i=1}^{n}\max(0,c^{i})^{2}\geq 0$,   i.e. encouraging positive components.
-(2) $\sum_{i=1}^{n}\min(0,c^{i})^{2}\leq 0$,  i.e. eliminating negative components.
-The absence of such constraints allows bias amplification when bias projection is present. Based on these findings, we establish the following:
-Theorem 1 (Condition for Bias Amplification).
-Bias amplification, a self-reinforcing process where a model trained on synthetic data amplifies the pre-existing biases from previous training, leading to an increased bias in the data it generates for future training, occurs if both bias projection and constraint deficiency are present.
-It occurs only if there is bias projection in at least one parameter dimension and constraint deficiency is present.
-Alongside understanding the causes of bias amplification, this also implies that bias amplification can occur even without model collapse.
-Specifically, statistical approximation and functional expressivity errors [^1] We have not discuss functional approximation error here, as it is a general concept that encompasses any structural setting, from learning to objective function, that may lead to suboptimal training outcomes, making it too general for this discussion.—common causes of model collapse (Shumailov et al., 2024)—are unrelated with bias projection and constraint deficiency.
-This is further supported by our experiments in Sections 3.2 and 5.
-3.2 Weighted Maximum Likelihood Estimation (WMLE)
-To demonstrate bias amplification in a controlled setting, we consider a statistical estimation cycle using WMLE, devoid of statistical approximation and functional expressivity errors.
-The mathematical formulation of WMLE is provided in Appendix A.
-Specifically, we generate a "biased" pre-training dataset $\mathcal{D}_{\text{pre}}$ consisting of $N=10,000$ samples from a "biased" Beta$(\alpha=3,\beta=2)$ distribution, with a mean of $x=0.6$.
-Maximum likelihood estimation (MLE) is used to estimate the probability density function (pdf) of this beta distribution, resulting in the ’biased’ pre-trained pdf (referred to as $f_{\text{pre}}$).
-Our goal in the estimation cycle is to estimate the pdf of the true distribution, Beta$(\alpha=2,\beta=2)$, with a mean of $x=0.5$.
-We generate $N$ samples from the true distribution, referred to as $D_{\text{real}}$, serving as data for the initial round of training.
-In the first round, we estimate the true pdf using WMLE with $D_{\text{real}}$ and weights derived from $f_{\text{pre}}$, simulating the impact of the biased pre-trained model during the current training.
-The newly estimated pdf is referred to as $f_{\text{Generation 0}}$.
-Then, we generate the synthetic data $D_{\text{Generation 0}}$ of size N using $f_{\text{Generation 0}}$.
-In the second round, we use WMLE with $D_{\text{Generation 0}}$ and weights derived from $f_{\text{Generation 0}}$ to obtain $f_{\text{Generation 1}}$.
-We then repeat this process, recursively estimating $f_{\text{Generation 2}}$ through $f_{\text{Generation 10}}$, simulating an iterative training process on synthetic data.
-Results:
-As depicted in Figure 1, the estimated distributions progressively shift towards the mean of the biased Beta$(\alpha=3,\beta=2)$ distribution, at $x=0.6$, becoming increasingly peaked with each generation.
-This behavior exemplifies bias amplification.
-The distortion arises from two sources: bias projection, where the estimation process emphasizes data regions where the ’biased’ pre-trained distribution assigns higher probability; and constraint deficiency, where WMLE lacks mechanisms to counteract the bias introduced by the weights, allowing the bias to be amplified over generations.
-For comparison, Figure 2 shows the results using standard MLE without weighting.
-The estimated distributions remain consistent across generations, accurately capturing the true Beta(2,2)22(2,2)( 2 , 2 ) distribution.
-This indicates that without the influence of biased pre-trained pdf and the weighted learning process, bias amplification does not occur.
-Figure 1: Weighted Maximum Likelihood Estimation (WMLE) over 10 generations.
-Figure 2: Standard Maximum Likelihood Estimation (MLE) over 10 generations, starting from the true distribution Beta(2,2)22(2,2)( 2 , 2 ) without weighting.
-4 Experiment Design
-In this section, we present and explain in detail the experiments designed to investigate the bias amplification of LLMs (see results in Section 5).
-Specifically, we analyze the sequential and synthetic fine-tuning of GPT-2.
-While similar experiments could be conducted with larger models such as GPT-4 or Claude in non-fine-tuning contexts, doing so would likely incur significant environmental costs.
-The focus of this study is on political bias within the US political spectrum, particularly in the context of sentence continuation tasks.
-This is important because LLMs are increasingly becoming a source of information for global news consumption Maslej et al.
+Intuitively, bias in the output arises from specific neurons, whose weights increase when their activations align with the fine-tuning dataset.
+If biased neurons capture more relevant patterns than unbiased ones, optimization will further reinforce their weights.
+The second is sampling error, akin to statistical approximation error Shumailov et al.
+(2024).
+If the model has a pre-existing bias, it inherently assigns higher probabilities to tokens that produce biased outputs.
+Consequently, during synthetic data generation, unbiased tokens—and thus unbiased samples—are more likely to be lost at each resampling step with a finite sample, though this error vanishes as the sample size approaches infinity.
+This overrepresents biased patterns in the synthetic data, surpassing the model’s original bias and true next-token probabilities.
+Sampling error thus complements bias projection by further activating biased neurons in response to the skewed dataset.
+By definition, bias projection is a sufficient condition for bias amplification, while sampling error serves as a complementary factor.
+However, sampling error is a sufficient condition for model collapse to occur with nonzero probability Shumailov et al.
+(2024).
+This distinction motivates our investigation into whether bias amplification can occur without model collapse, as this would necessitate additional mitigation strategies beyond those for model collapse.
+3.2 Statistical Simulation
+To simulate a controlled setting without sampling error, we consider a statistical estimation cycle using WMLE [^1] The mathematical formulation is detailed in Appendix A.
+with a large sample size of each resampling step.
+Specifically, we generate a pre-training dataset $\mathcal{D}_{\text{pre}}$ with 100,000 samples from a Beta$(3,2)$ distribution, representing a biased pre-training dataset.
+Using maximum likelihood estimation (MLE), we estimate its probability density function, yielding the pre-trained model $f_{\text{pre}}$.
+Next, we fine-tune $f_{\text{pre}}$ to approximate a different distribution, Beta(2,2)22(2,2)( 2 , 2 ).
+We generate 100,000 samples from this distribution, denoted as $D_{\text{real}}$, which serves as the initial fine-tuning dataset.
+In the first round, we apply weighted maximum likelihood estimation (WMLE) using weights derived from $f_{\text{pre}}$, which encode the pre-existing bias of the pre-trained model.
+This weighting captures the influence of the pre-trained model’s parameters on subsequent training.
+This produces the fine-tuned model $f_{\text{0}}$.
+We then generate a synthetic dataset $D_{\text{0}}$ of the same size using $f_{\text{0}}$, initiating the iterative fine-tuning loop.
+In each subsequent round, WMLE is applied using $D_{\text{k}}$ with weights from $f_{\text{k}}$, resulting in $f_{\text{k+1}}$.
+This process is repeated iteratively, producing models $f_{\text{1}}$ through $f_{\text{10}}$.
+Figure 2(a) shows the estimated distributions gradually shift toward the mean of the biased pre-training dataset at $x=0.6$, becoming progressively more peaked over generations.
+This occurs despite further training on samples drawn from Beta(2,2)22(2,2)( 2 , 2 ) and synthetic data generated from successive models.
+The distortion arises because the fine-tuning process disproportionately emphasizes regions where the pre-trained distribution assigns higher probability, leading to biased learning.
+For comparison, Figure 2(b) presents the results using standard MLE without weighting.
+In this case, the estimated distributions remain stable across generations, accurately representing the Beta(2,2)22(2,2)( 2 , 2 ) distribution.
+4 Experimental Design
+This section provides the details of the experiments on LLMs, focusing on the sequential and synthetic fine-tuning of GPT-2 under different setups [^2] Similar experiments with larger models like GPT-4 or Claude can be easily adapted to the current setup but would incur significant environmental costs..
+The step-by-step experimental procedure is outlined in Figure 1.
+Our study focuses on the political bias of LLMs within the US political spectrum, particularly in sentence continuation tasks.
+This is important as LLMs are increasingly influencing global news consumption Maslej et al.
 (2024); Peña-Fernández et al.
-(2023); Porlezza and Ferri (2022), and traditional news outlets, such as the Associated Press, are beginning to incorporate LLMs to automate content generation from structured data (The Associated Press, 2024).
-4.1 Experiment Dataset Preparation
+(2023); Porlezza and Ferri (2022), and traditional news outlets, such as the Associated Press, are beginning to integrate LLMs for automated content generation from structured data (The Associated Press, 2024).
+4.1 Dataset Preparation
 We randomly selected 1,518 articles from the Webis-Bias-Flipper-18 dataset (Chen et al., 2018), which contains political articles from a range of U.S.
 media outlets published between 2012 and 2018, along with bias ratings assigned at the time for each media source.
-These ratings were provided by Allsides and were determined through a multi-stage process involving both bipartisan experts and members of the general public (AllSides, 2024a).
-Based on the bias ratings, the random selection was stratified to ensure that the 1,518 articles were evenly divided into three groups of 506 articles each, representing left-leaning, right-leaning, and center-leaning media.
-4.2 Fine-tuning and Synthetic Data Generation
-Inspired by Shumailov et al.
+These bias ratings, provided by AllSides, were determined through a multi-stage process incorporating assessments from both bipartisan experts and the general public (AllSides, 2024a).
+The random sampling was stratified based on bias ratings to ensure an even distribution of the 1,518 articles into three groups of 506 each, representing left-leaning, right-leaning, and center-leaning media.
+4.2 Successive Fine-tuning
+Building on the approach in Shumailov et al.
 (2024); Dohmatob et al.
-(2024b), each training cycle begins with GPT-2 fine-tuned on the dataset of human-written political articles.
-The fine-tuned GPT-2 is then used to generate the first set of synthetic data, which is subsequently employed to further fine-tune the model, producing model generation 1.
+(2024b), each training cycle begins with fine-tuning GPT-2 on a dataset of 1,518 news articles.
+The fine-tuned model, referred to as Generation 0, generates a synthetic dataset $D_{0}$ of the same size, which is then used for fine-tuning to produce Generation 1.
 This iterative process continues until Generation 10 is reached.
-For details on the fine-tuning setup, refer to Appendix B.
-[…]
+Note that synthetic fine-tuning starts at Generation 1.
+4.3 Synthetic Data Generation
+Synthetic datasets, $\{D_{i}\}_{i=0}^{10}$, are generated as follows: Each original news article is tokenized into 64-token blocks, and for each block, the model predicts the next 64 tokens using deterministic generation to enhance the replicability of the study.
+The generated tokens are then decoded back into text, producing a synthetic dataset of the same size as the original.
 4.4 Political Bias Metric
-We develop a metric model to classify the political leaning of synthetic outputs generated by large language models.
-The classifier assigns labels corresponding to left, center, and right-leaning, trained on data from AllSides (AllSides, 2024a), a source that assigns bias ratings to each media outlet.
-The methodology for the classifier is as follows:
-Dataset and Preprocessing.
-We use a dataset consisting of political articles with labeled bias categories for the source media: left, center, and right.
-We concatenate the article titles with their body text to form the input.
-To address class imbalance, we resample the center-leaning articles, balancing the dataset with respect to each category.
-The dataset is then split into training (70%), validation (15%), and test (15%) subsets, stratified by bias label.
-Moreover, we conduct a human review of the test subsets to eliminate any identifiable information about media sources or writers.
-Model Architecture.
-We experiment with multiple transformer-based models, including BERT (Devlin et al., 2018), RoBERTa (Liu et al., 2019), and other state-of-the-art architectures, to select the best-performing model based on macro F1 score.
-Each model is fine-tuned using the HuggingFace ‘Trainer‘ class, with a learning rate of $2\times 10^{-5}$, a batch size of 16, and 5 training epochs.
-We employ a cross-entropy loss function for multi-class classification.
-Training Procedure.
-Tokenization is performed using each model’s respective tokenizer with a maximum sequence length of 512 tokens.
-To mitigate overfitting, weight decay of 0.01 is applied during training.
-The model checkpoints are saved after each epoch, and the best model is selected based on macro F1 score evaluated on the validation set.
-We use a weighted random sampler during training to ensure balanced class representation.
-Evaluation Metrics.
-We evaluate the models using the macro F1 score to account for the multi-class nature of the task, ensuring that performance is balanced across all bias categories.
-The final evaluation is conducted on the held-out test set.
-Additionally, we report the loss, runtime, and sample processing rates for completeness.
-Results.
-After conducting a grid search across several models, we find that the roberta-base model achieves the best performance with an evaluation loss of 0.4035 and a macro F1 score of 0.9196 on the test set (see Table 1).
-Thus, roberta-base model is selected as the benchmark for political bias detection in subsequent experiments.
-Table 1: Evaluation Results for Political Leaning Classifier
+We develop a classification model to assess the political leaning of each LLM based on its generated synthetic news articles.
+The model is trained on the Webis-Bias-Flipper-18 dataset, excluding the 1,518 articles used for GPT-2 fine-tuning.
+To mitigate class imbalance, center-leaning articles are resampled to ensure equal representation across categories.
+The dataset is then divided into training (70%), validation (15%), and test (15%) subsets, stratified by bias label.
+Additionally, we conduct a human review to remove any identifiable information about media sources and authors.
+After performing a grid search across multiple models, we find that roberta-base achieves the best performance, with an evaluation loss of 0.4035 and a macro F1 score of 0.9196 on the test set (see Table 1).
+Thus, we select roberta-base as the benchmark for political bias detection in subsequent experiments.
+Details on model training are provided in Appendix C.
+Table 1: Evaluation Results for Political Bias Classifier
 Model Macro F1 Score
@@
 4.5 Generation Quality Metric
-We introduced a metric for evaluating generation quality, specifically addressing the issue of repetitive content in later model iterations, which can skew traditional perplexity metrics.
+We introduce a metric to evaluate generation quality, specifically addressing the issue of repetitive content in later model iterations, which can distort traditional perplexity metrics.
 This metric is based on the Gibberish Detector (Jindal, 2021), which identifies incoherent or nonsensical text.
-The detector classifies text into four categories: (1) Noise, where individual words hold no meaning; (2) Word Salad, where phrases are incoherent; (3) Mild Gibberish, where grammatical or syntactical errors distort meaning; and (4) Clean, representing coherent, meaningful sentences.
-To evaluate generation quality, we assign the Gibberish score to each sentence: 3 for Clean, 2 for Mild Gibberish, 1 for Word Salad, and 0 for Noise.
-An average score is calculated for each article, resulting in the text quality index.
-It focuses on the coherence and meaning of the generated text and outperforms perplexity in accurately reflecting actual generation quality, as demonstrated in Section 5.2.
+The detector categorizes text into four levels: (1) Noise—individual words hold no meaning, (2) Word Salad—incoherent phrases, (3) Mild Gibberish—grammatical or syntactical distortions, and (4) Clean—coherent, meaningful sentences.
+To quantify generation quality, each sentence receives a Gibberish score: 3 for Clean, 2 for Mild Gibberish, 1 for Word Salad, and 0 for Noise.
+The text quality index is computed as the average score across all sentences in an article.
+This metric prioritizes coherence and meaning, offering a better assessment of generation quality than perplexity.
 5 Results
-In this section, we present the results on bias and generation performance of GPT-2 throughout iterative training, both with and without the deployment of mitigation strategies.
+In this section, we analyze the evolution of political bias and generation quality in GPT-2 over successive iterations of synthetic fine-tuning, comparing results with and without mitigation strategies.
 5.1 Political Bias
-Figure 3: This figure displays the distribution of articles across the political bias labels—’Left’, ’Center’, and ’Right’—for GPT-2’s outputs, based on classifications made by the classifier (see Section 4.3).
-Using an unbiased dataset of 1,518 human-written articles evenly distributed across three political-leaning categories, we generated 1,518 synthetic articles by deterministically predicting the next 64 tokens based on the previous 64 tokens.
-If GPT-2 had no pre-existing bias, the resulting distribution would be evenly spread across the three political-leaning labels.
-However, as shown in Figure 3, the model generates more center (47.9%) and right-leaning articles (46.8%), indicating a bias toward these categories prior to any fine-tuning.
-Figure 4: This figure shows the percentage of right-leaning articles across multiple generations for three generation setups, with ’Synthetic’ representing the deterministic setup.
-We then fine-tuned GPT-2 to produce Generation 0 and continued the process iteratively using synthetic datasets, up to Generation 10.
-As shown in Figure 4, the fine-tuned model displayed a significant amplification of bias toward the Right across all three generation setups after Generation 2.
-Interestingly, Generation 0 in the deterministic setup (the ’Synthetic’ line), which was fine-tuned on an unbiased dataset of real articles, produces an even higher percentage of right-leaning articles (54.5%) compared to the unfine-tuned GPT-2.
-To ensure bias classification aligns with human intuition, we further examine model responses with human assistance.
-For illustration, we analyze a political opinion article, titled "First Read: Why It’s So Hard for Trump to Retreat on Immigration", from NBC News, a left-leaning outlet rated by AllSides (NBC News, 2016; AllSides, 2024b), critiquing President Trump’s immigration policies.
-Using a qualitative analysis grounded in foundational media studies (Entman, 1993; Rodrigo-Ginés et al., 2024; Groeling, 2013), we observe a significant rightward shift between generations 0 and 4 for this example, consistent with the classifier’s results (see Appendix C).
-As the generations progress, the synthetic texts increasingly frame Trump’s immigration policies as strong and effective.
-While the original article focuses on the dilemma and electoral interests behind Trump’s stance, Generation 0 begins to emphasize his determination and reliability, omitting the criticism present in the original.
-By Generation 4, the narrative shifts further, focusing almost entirely on portraying Trump’s personal qualities and electoral legitimacy, with statements like "he is not a politician, he is a man of action." Notably, from Generation 0 onward, the term "undocumented immigrant" in the original article is replaced with "illegal immigrants."
+Figure 3: This figure displays the distribution of articles across the political bias labels—’Left’, ’Center’, and ’Right’—for GPT-2’s outputs, based on classifications made by Political Bias Metric.
+GPT-2 was used to generate the synthetic dataset.
+Since the original human-written dataset is unbiased—with an equal number of articles for each political-leaning category—the synthetic dataset should ideally mirror this balanced distribution if GPT-2 had no pre-existing bias.
+Figure 3 presents the distribution of synthetic articles generated by GPT-2 across political bias labels.
+The model predominantly produces center-leaning (47.9%) and right-leaning (46.8%) articles, suggesting a pre-existing bias for these categories before any fine-tuning.
+Starting from the initial GPT-2 model, we fine-tuned it iteratively, generating synthetic datasets to train successive models up to Generation 10.
+Figure 4(a) illustrates how bias amplifies across generations.
+Surprisingly, fine-tuning on the unbiased real dataset increases right-leaning bias, with 53.7% of articles classified as right-leaning in Generation 0.
+Furthermore, without mitigation strategies, successive rounds of synthetic fine-tuning lead to a continuous rise in right-leaning articles, peaking at Generation 6 (67.6%) before
+(a) Bias performance
+(b) Generation quality
+Figure 4: Evolution of bias performance and generation quality across generations, with Generation 0 representing GPT-2 fine-tuned on the unbiased dataset.
+The baseline (’Synthetic’) is compared with three mitigation strategies, and the text quality index is shown with 95% confidence intervals.
+stabilizing.
+Figures 7 and 8 in Appendix D show the percentage of center-leaning and left-leaning articles across generations.
+Notably, the proportion of of center-leaning articles remains stable at approximately 35% throughout synthetic fine-tuning.
+To further illustrate, we analyze how a specific article, "First Read: Why It’s So Hard for Trump to Retreat on Immigration", evolves in the synthetic generations.
+This case study reveals a progressive rightward shift in framing and word choice, mirroring the classifier’s results (Appendix E).
+As generations progress, the synthetic texts increasingly depict Trump’s immigration policies as strong and effective.
+While the original article highlights the dilemmas and electoral considerations behind Trump’s stance, Generation 0 begins to emphasize his determination and reliability, omitting the critical perspectives present in the original.
+By Generation 4, the narrative shifts even further, focusing almost entirely on portraying Trump’s personal qualities and electoral legitimacy, with statements such as "he is not a politician, he is a man of action." Notably, starting from Generation 0, the term "undocumented immigrant" in the original article is consistently replaced with "illegal immigrants."
 5.2 Generation Quality
-Figure 5: This figure shows the average text quality index of synthetic articles, along with 95% confidence intervals, across generations for three generation setups, with ’Synthetic’ representing the deterministic setup.
-As shown in Figure 5, there is a model collapse in terms of the deterioration of the generation quality across all three generation setups: the average text quality index (defined in Section 4.4) decreases over the iterative training.
-Moreover, the distributions of the text quality index for all three generation setups shift markedly toward the low-quality region and starts to produce data that is never produced by Generation 0, as shown by the distribution plots in Appendix D.
-These findings are consistent with the existing literature on model collapse, such as (Shumailov et al., 2024).
-However, we did not observe significant variation in variance, as seen in Figure 5 and distribution plots in Appendix D.
-In contrast, perplexity measurements show a steady decrease across generations for the deterministic and beam search generation setups, generally indicating improved generation quality (see Figure 11 in Appendix E).
-For a closer look, as shown by the examples in Appendix F, the generated articles progressively lose coherence and relevance with each successive generation, as repetition and fragmented sentences become more common.
-By Generation 10, the text is largely incoherent and disconnected from the original content, diminishing its readability and meaning.
-However, despite the clear deterioration in generation quality, perplexity decreases over generations, as shown by the results at the end of each synthetic output example.
-This pattern is observed in most other synthetic outputs as well, suggesting that perplexity fails to accurately reflect the model’s true generation capabilities and may be biased toward repetitive content.
+Figure 4(b) illustrates the text quality index across generations.
+In the training loop without any mitigation strategy, model collapse occurs, as evidenced by the gradual decline in the average text quality index.
+Furthermore, the distribution of the text quality index shifts significantly toward the lower-quality region over generations, eventually generating data that was never produced by Generation 0 (Figure 9 in Appendix F).
+These results align with prior research on model collapse, such as (Shumailov et al., 2024), though we did not observe substantial variation in variance.
+Conversely, perplexity measurements exhibit a consistent decline across generations, generally suggesting an improvement in generation quality (Figure 10 in Appendix G).
+For a closer look, the examples in Appendix H illustrate how generated articles gradually lose coherence and relevance across generations, with increasing occurrences of repetition and fragmented sentences.
+By Generation 10, the text becomes largely incoherent and detached from the original content, reducing its readability and meaning.
+However, despite the evident decline in generation quality, perplexity decreases over generations, as indicated by the results at the end of each synthetic output example.
+This pattern is consistent across most synthetic outputs, suggesting that perplexity does not accurately capture the model’s true generative capabilities and is prone to artificial inflation in the presence of frequent repetitions.
 5.3 Mitigation Strategies
-Figure 6: This figure shows the percentage of right-leaning articles across multiple generations for three mitigation strategies, with ’Synthetic’ representing the baseline deterministic setup.
-Figure 7: This figure shows the average text quality index of synthetic articles, along with 95% confidence intervals, across multiple generations for three mitigation strategies, with ’Synthetic’ representing the baseline deterministic setup.
-We applied three potential mitigation strategies under the deterministic generation setup: (1) Overfitting, which involved increasing the training epochs to 25 (five times the baseline) and setting weight decay to 0 to reduce regularization and encourage overfitting, as proposed by Taori and Hashimoto (2022) based on their uniformly faithful theorem of bias amplification; (2) Preserving 10% of randomly selected real articles during each round of synthetic fine-tuning, a method proposed and used in (Shumailov et al., 2024; Alemohammad et al., 2023; Dohmatob et al., 2024b; Guo et al., 2024); and (3) Accumulating all previous fine-tuning datasets along with the new synthetic dataset in each fine-tuning cycle, which was introduced by Gerstgrasser et al.
+We applied three mitigation strategies: (1) Overfitting, which involved increasing the training epochs to 25 (five times the baseline) and setting weight decay to 0 to reduce regularization and encourage overfitting, as proposed by Taori and Hashimoto (2022) based on the uniformly faithful theorem of bias amplification; (2) Preserving 10% of randomly selected real articles during each round of synthetic fine-tuning, a method proposed and used in
+(a) Neuron weights vs. bias performance.
+(b) Neuron weights vs. generation quality.
+Figure 5: Scatter plots showing the Pearson correlation between neuron weights and model behavior.
+(Shumailov et al., 2024; Alemohammad et al., 2023; Dohmatob et al., 2024b; Guo et al., 2024); and (3) Accumulating all previous fine-tuning datasets along with the new synthetic dataset in each fine-tuning cycle, which was introduced by Gerstgrasser et al.
 (2024).
-As shown in Figure 6, overfitting helps reduce bias amplification in the early generations compared to the no-mitigation setup (the ’Synthetic’ line), but it fails to prevent bias amplification in the later generations.
-Additionally, it incurs a significant cost—further deterioration in generation quality, as shown in Figure 7.
-Notably, both the preservation and accumulation strategies perform well in mitigating both bias amplification and model collapse.
-Interestingly, these strategies even result in bias reduction.
+As shown in Figure 4(a), overfitting helps reduce bias amplification in the early generations compared to the no-mitigation setup (the ’Synthetic’ line), but it fails to prevent bias amplification in the later generations.
+Additionally, it incurs a significant cost—further deterioration in generation quality, as shown in Figure 4(b).
+Notably, both the preservation and accumulation strategies effectively mitigate model collapse and reduce bias, yielding 41.89% and 42.7% right-leaning articles, respectively, at Generation 10.
 5.4 Mechanistic Interpretation
-To gain a clearer understanding of the causes of bias amplification and how it empirically differs from model collapse, we investigate how neurons behave and vary across different versions  [^2] We have 11 generations for each training setup and a total of 6 setups, resulting in 66 versions of fine-tuned GPT-2.
+To gain a clearer understanding of the causes of bias amplification and how it empirically differs from model collapse, we investigate mechanistically how neurons behave and vary across different generations  [^3] We have 11 generations for each training setup and a total of 6 setups, resulting in 66 versions of fine-tuned GPT-2.
 of fine-tuned GPT-2 with differing levels of generation quality and bias performance.
 First, we examined how the weight of each neuron changes across different versions.
-For each of the 9,216 neurons, we calculated the correlation between its weight and the model’s bias performance (i.e.
-the percentage of right-leaning articles) across the 66 versions, as shown in Figure 12 in Appendix G.
-To statistically test the significance of the correlations, we calculated Newey-West adjusted p-values for the 9,216 neurons, using the null hypothesis that no significant relationship exists between the change in neuron weight and the change in bias performance, and applying the Bonferroni correction to maintain an overall 0.05 significance level.
-This yielded a set of 553 neurons whose weight value changes are significantly correlated with changes in bias performance (i.e.
-$p\text{-value}<1.36\times 10^{-6}$), indicating they are likely primary contributors to the shift in bias  [^3] All test results mentioned in this section are uploaded.
-In contrast, we identified only one neuron whose weight changes are significantly correlated with changes in generation quality.
-This suggests that shifts in bias performance and generation quality are likely driven by different underlying mechanisms.
-The formula for the statistical tests is provided in Appendix K.
-Secondly, we examined how the activation values of each neuron change across different versions of GPT-2 when provided with the same input used to generate synthetic data (see Section 4.2).
-Specifically, we attached activation hooks to capture activation values (i.e., outputs from these layers during a forward pass) for 9,216 neurons.
-With the extracted activation values, we identified two sets: one consisting of 3,062 neurons whose activation value changes are significantly correlated with bias performance changes (i.e.
-$p\text{-value}<1.36\times 10^{-6}$), and another consisting of only 2 neurons whose activation value changes are significantly correlated with changes in generation quality.
-6 Discussion
-We now discuss the implications of bias amplification for the future development of large language models.
-As demonstrated, bias amplification is a parallel concept to model collapse and warrants equal emphasis in future research.
-In our case, we successfully identified ad hoc mitigation strategies, such as Preservation and Accumulation, which act as natural constraints on gradient updates during training by ensuring the model still considers the initial unbiased dataset.
-However, if the training dataset itself is biased or overrepresents certain demographic or political groups, these methods may fail, while pre-processing strategies could be too resource-intensive.
-Therefore, there is an urgent need to develop in-process mitigation strategies to address the constraint deficiency and promote more equitable and fair model development.
+For each of the 9,216 neurons, we calculated the correlation between its weight and the model’s bias performance (or generation quality) across the 66 versions, as shown in Figure 5.
+To statistically test the significance of the correlations, we estimate the linear model:
+$\Delta y_{i}=\alpha_{j}+\beta_{j}\Delta x_{i,j}+\epsilon_{i,j}$ (2)
+where $\Delta y_{i}$ represents the change in the proportion of right-leaning articles (or text quality index) generated between model $i-1$ and model $i$, and $\Delta x_{i,j}$ denotes the corresponding change in the weight of neuron $j$ during the same transition.
+The coefficient $\beta_{j}$ captures the degree to which changes in the weight of neuron $j$ influence shifts in political bias (or generation quality), while $\alpha_{j}$ serves as a constant term, and $\epsilon_{i,j}$ represents the residual error.
+By applying a first-order difference to both $x_{i,j}$ and $y_{i}$, we mitigate potential serial correlations, ensuring that our regression estimates more accurately reflect the dynamic impact of individual neuron weight updates on bias amplification.
+[…]
+5.5 Further Investigation
+(a) Bias performance
+(b) Generation Quality
+Figure 6: Evolution of bias performance and generation quality across generations, with Generation 0 representing GPT-2 fine-tuned exclusively on center-leaning news articles.
+The baseline (’Synthetic’) is compared with three mitigation strategies, and the text quality index is shown with 95% confidence intervals.
+We conduct an alternative synthetic training cycle, beginning with GPT-2 fine-tuned on 1,518 randomly sampled center-labeled articles.
+We compare the baseline setup with the most effective and cost-efficient mitigation strategy identified in our previous results: Preservation.
+As shown in Figure 6, Preservation successfully prevents model collapse but fails to mitigate bias amplification in center-leaning article generation, which increases from 72.9% at Generation 0 to 88.2% at Generation 10.
+This result suggests that while reducing sampling error through techniques like Preservation effectively mitigates model collapse, it does not necessarily prevent bias amplification, thereby validating our mechanistic interpretation.
+6 Discussion and Conclusion
+We now explore the implications of our findings for the future research on large language models.
+Our results demonstrate that bias amplification operates through a distinct underlying mechanism from model collapse, as supported by both theoretical intuition and empirical evidence.
+Theoretically, the primary driven factor of model collapse is sampling error, but sampling error is only a auxiliary factor in bias amplification which means it is not a necessary or sufficient condition.
+Therefore, the mitigation strategy targeting sampling error is not necessarily helping with mitigating bias amplification.
+Empirically, we found mitigation strategies like preservation very effective in mitigating model collapse but failed at bias amplification in some cases.
+Even in cases that it helps with both, we do identify a distinct set of neurons responsible for the two phenomenons.
+Intuitively, the main reason for them to work on model collapse is, the preservation and accumulation propose a natural constraint on the learning process by recalling the real dataset in further synthetic training.
+However, when the real dataset itself is biased, the recalling behavior only raise up the dominance of biased patterns in the further training dataset.
+Indeed, applying bias-category-weighted sampling in preservation or accumulation strategies may help mitigate bias amplification.
+However, this approach inherently introduces additional sampling error, which could, in turn, incur model collapse.
+Thereby, This highlights the urgent need for more targeted and efficient mitigation strategies specifically addressing bias amplification to ensure fairer and more equitable model development.
+To develop such targeted mitigation strategies, a deeper mechanistic understanding of bias amplification is essential.
+In our analysis, we adopt a statistical approach rather than Sparse Autoencoder (SAE) methods due to our focus on tracking the temporal dynamics of bias amplification across generations.
+This approach allows us to examine how neuron weights evolve over iterations and how these changes correlate with model bias, whereas existing SAE pipelines are primarily suited for static analysis.
+Additionally, political bias is a more nuanced concept compared to harmful or discriminatory outputs.
+It is characterized by the disproportionate overrepresentation or beautification of a particular political leaning’s ideas in a model’s generation.
+If content from different political perspectives is generated in a balanced manner, the model is not politically biased.
+Given this definition, pinpointing neurons responsible for such disproportionality using SAE is particularly challenging.
+Future research could focus on refining mechanistic analysis techniques for political bias and uncovering more effective ways to constrain bias amplification during synthetic model training.
 7 Limitations
@@
 media.
-Additionally, our experiments were conducted using GPT-2, a relatively smaller model compared to state-of-the-art architectures like GPT-4 or LLaMA-2.
-Future research should extend our empirical approach to other contexts and larger LLMs.
+The political spectrum may shift over time, necessitating periodic updates to the political bias classifier to ensure its accuracy when benchmarking recent datasets.
+Additionally, due to resource constraints, our experiments were conducted using GPT-2, a relatively small language model.
+Future work may extend our methodology to models with larger architectures.
+However, it’s important to note that employing a larger model, such as DeepSeek-V3, which relies on a considerably larger pre-training dataset and has many more parameters, will likely require larger synthetic fine-tuning datasets to effectively demonstrate its impact.
+For instance, given that our current fine-tuning datasets average 777,216 tokens, transitioning from GPT-2 to DeepSeek-V3 could necessitate much larger datasets and incur significant fine-tuning costs.
 Another limitation lies in our choice of mitigation strategies.
-While Preservation and Accumulation show promise in reducing bias amplification, their computational cost and scalability must be considered.
-Moreover, the mitigation strategies were tested primarily in the context of synthetic data generation, and their efficacy in real-world deployments requires further investigation.
+While Preservation and Accumulation show promise in reducing model collapse, their computational cost and scalability must be considered.
+Moreover, the mitigation strategies were tested primarily in the context of synthetic fine-tuning, and their efficacy in real-world deployments requires further investigation.
 8 Ethical Considerations
-This study addresses bias amplification in LLMs, a technical phenomenon with profound ethical implications, particularly regarding fairness and the integrity of AI systems.
+This study focus on bias amplification in LLMs, a phenomenon with profound ethical implications, particularly regarding fairness and the integrity of AI systems.
 The risk of bias amplification is especially concerning in systems that are iteratively trained on synthetic data, as it can lead to unintended distortions in model outputs.
@@
 This amplification can mislead users or result in models perpetuating one-sided perspectives, which could be especially problematic in sensitive domains like news summarization, policy generation, or social media content moderation.
-Moreover, the identification of distinct neural mechanisms for bias amplification and model collapse brings to light the challenges of ensuring equitable performance across all dimensions of model behavior and raises important ethical questions regarding the adequacy of current mitigation techniques, particularly in high-stakes scenarios where the cost of algorithmic bias is substantial.
-Future research should prioritize the development of more comprehensive and domain-specific bias mitigation techniques, with a clear focus on minimizing ethical risks.
-Additionally, rigorous testing and validation across diverse datasets and real-world applications will be critical to ensuring that models trained using these methods do not exacerbate existing inequalities or produce harmful outcomes.
+It is crucial to explicitly state that the methodologies and data used in this research should not be applied to develop or train biased models for harmful applications.
+This study aims to enhance the understanding of bias amplification and model collapse in LLMs while promoting responsible and ethical AI development.
+This work includes content that may contain personally identifying information or offensive language.
+However, all such material is derived solely from publicly available news article datasets or generated synthetically by models fine-tuned on these open-source datasets—or on synthetic data produced by earlier generations in our training cycle.
+Consequently, any sensitive or offensive content reflects the characteristics of the source material rather than our endorsement of it.
+Our objective is to investigate and understand political bias in LLMs so that strategies can be developed to prevent such content from disproportionately appearing in real-world deployments.
+Additionally, we conduct a human review of the news article dataset to remove any identifiable information about article authors.
 Appendix A Mathematical Formulation of WMLE
@@
 In Maximum Likelihood Estimation, we maximize the log-likelihood with $w_{j}=1,\forall j$.
-Appendix B Fine-tuning Setup
+Appendix B Details on Fine-tuning for LLMs
 The fine-tuning setup remained consistent across all experiments unless stated otherwise: the input length was capped at 512 tokens with the EOS token used for padding.
@@
 Fine-tuning was conducted using the Hugging Face ‘Trainer‘ class, and after each cycle, the model was saved for generating synthetic data for the subsequent iteration.
-Appendix C Qualitative Bias Analysis Framework and Example of Bias Amplification Across Generations
+Appendix C Details on Model Training for Political Bias Metric
+We experiment with multiple transformer-based models, e.g.
+BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019), to select the best-performing model based on macro F1 score.
+Each model is fine-tuned using the HuggingFace ‘Trainer‘ class, with a learning rate of $2\times 10^{-5}$, a batch size of 16, and 5 training epochs.
+We employ a cross-entropy loss function for multi-class classification.
+Tokenization is performed using each model’s respective tokenizer with a maximum sequence length of 512 tokens.
+To mitigate overfitting, weight decay of 0.01 is applied during training.
+The model checkpoints are saved after each epoch, and the best model is selected based on macro F1 score evaluated on the validation set.
+We use a weighted random sampler during training to ensure balanced class representation.
+We evaluate the models using the macro F1 score to account for the multi-class nature of the task, ensuring that performance is balanced across all bias categories.
+The final evaluation is conducted on the held-out test set.
+Additionally, we report the loss, runtime, and sample processing rates for completeness.
+Appendix D Percentage of Center (Left) Biased Articles
+Figure 7: This figure shows the percentage of center-leaning articles across multiple generations, comparing the baseline setup (’Synthetic’) with three mitigation strategies.
+Figure 8: This figure shows the percentage of left-leaning articles across multiple generations, comparing the baseline setup (’Synthetic’) with three mitigation strategies.
+Appendix E Qualitative Bias Analysis Framework and Example of Bias Amplification Across Generations
 We employed qualitative methods to confirm our findings in media bias.
 Specifically, we utilized a media bias identification framework grounded in foundational works such as Entman’s framing theory (Entman, 1993) and other research on media bias detection (Rodrigo-Ginés et al., 2024; Groeling, 2013).
-This framework provides a robust lens to evaluate political biases in the framing and language use of media texts (explained below).
+This framework provides a robust lens to evaluate political biases in the framing and language use of media texts.
 Given the nature of our data—text exclusive of visual or contextual cues like formatting—certain types of media bias commonly seen in formatted articles or televised programs (e.g., visual bias or tone) may not apply.
@@
 For example, "Trump can’t ignore a base that has cheered his uncompromising immigration position" and "And then there are the Trump supporters who’ve cheered the GOP presidential nominee for being so uncompromising on immigration." Additionally, "But here’s why it’s so hard — if not impossible — for Trump to retreat on immigration: He’s caught between his clear, unambiguous past statements and a base that might not be willing to see him moderate on the issue."
-    Analysis of Loaded Language Bia:
+    Analysis of Loaded Language Bias:
     Differences in Connotation.
@@
 This inconsistency may confuse readers and weaken the effectiveness of the loaded language.
-Appendix D Distribution of Text Quality Index Across Generations
-D.1 Deterministic Setup
-Figure 8: This figure presents the distribution of the text quality index across multiple generations of synthetic articles, for the deterministic setup of GPT-2.
-D.2 Beam Search Setup
-Figure 9: This figure presents the distribution of text quality index across multiple generations of synthetic articles, for the beam search setup of GPT-2.
-D.3 Nucleus Sampling Setup
-Figure 10: This figure presents the distribution of text quality index across multiple generations of synthetic articles, for the nucleus sampling setup of GPT-2.
-Appendix E Average Perplexity Across Generations
-Figure 11: This figure shows the average perplexity for synthetic articles , along with 95% confidence intervals, across multiple generations for three generation setups, with ’synthetic’ representing the determistic setup
-Appendix F Example of Quality Deterioration Across Generations
+Appendix F Distribution of Text Quality Index
+Figure 9: This figure presents the distribution of the text quality index across multiple generations for the baseline setup.
+Appendix G Average Perplexity Across Generations
+Figure 10: This figure shows the average perplexity for synthetic articles , along with 95% confidence intervals, across multiple generations for the baseline setup.
+Appendix H Example of Quality Deterioration Across Generations
 Examples of GPT-2 text outputs affected by iterative synthetic (Example articles are truncated for brevity, as the full articles are lengthy but exhibit consistent quality degradation throughout; see the dataset for details).
-F.1 Example of Deterministic Setup
+H.1 Example of Deterministic Setup
   • Original Article: The world’s eyes are on President Obama this week as he rallies a coalition of nations to “destroy” the extremist Islamic State (IS) and its terrorist-led territory.
@@
 Text Quality Index: 2.58; Perplexity: 6.68
-  • Generation 1: , says US official, says US military is ready to go, says Pentagon official.
-Says US military is ready to go, says Pentagon official.
-Says US military is ready to go, says Pentagon official.
-Says US military is ready to go, says Pentagon official.
-Says US military is ready to go, says Pentagon official the world is not so simple.
-Text Quality Index: 2.44; Perplexity: 5.77
-  • Generation 5: in Iraq and Syria (ISIS) group Read more The Iraqi army has been fighting the Islamic State since the group seized large swaths of territory in Iraq and Syria (ISIS) group in 2014.
+  • Generation 4: in Iraq and Syria (ISIS) group Read more The Iraqi army has been fighting the Islamic State since the group seized large swaths of territory in Iraq and Syria (ISIS) group in 2014.
 The Iraqi army has been fighting the Islamic State the Iraqi army.
@@
 Text Quality Index: 1.24; Perplexity: 4.23
-F.2 Example of Beam Search Setup
-  • Original Article: Senate approval will be a much higher bar to clear: There’s no shortage of bipartisan criticism from the upper chamber right now, and the GOP holds a considerably smaller majority margin on that half of the Hill.
-Text Quality Index: 2.70
-  • Generation 0: on Thursday, with Senate Majority Leader Mitch McConnell (R-Ky.) announcing that the bill would be headed to the House for a vote.
-Senate Majority Leader Harry Reid (D-Nev.) said in a statement that the Senate will vote on the bill ""as soon as possible." Text Quality Index: 2.26; Perplexity: 7.99
-  • Generation 1: Read more “It’s going to take a lot of work,” House Speaker John Boehner (R-Ohio) said on ABC’s “This Week.” “It’s going to take a lot of work,” House Speaker John Boehner (R-Ohio.
-“It’s going to take a lot of work,” said Sen.
-John McCain, R-Ariz., chairman of the Senate Armed Services Committee.
-“It’s going to take a lot of patience,” he added.
-“It’s going to take Recovery and Reinvestment Act of 2009.Text Quality Index: 2.26; Perplexity: 5.42
-  • Generation 5: committee MORE (R-Ohio) and Senate Majority Leader Mitch McConnell Addison (Mitch) Mitchell McConnellSchumer: Nunes gave Trump’secretly altered’ version of memo Davis: ‘Deep state’ existed in “ House Speaker John A.
-Boehner House Speaker House Speaker House Speaker John A.
-Boehner of the House Speaker House Speaker House Speaker John A.
-Boehner (R-Ohio) and Senate Majority Leader Mitch McConnell Addison (Mitch) Mitchell McConnellSchumer: Nunes gave Trump’secretly altered’ version of memo Davis: ‘Deep state’ existed in ’16 – but it elected Trump Health Care’s “I don’t know.
-Text Quality Index: 1.07; Perplexity: 3.33
-  • Generation 10: vote House Speaker John A.
-Boehner (R-Ohio) and Senate Majority Leader Mitch McConnell Addison (Mitch) Mitchell McConnellSchumer: Nunes gave Trump’secretly “ House Speaker John A.
-Boehner (R-Ohio) House Speaker House Speaker House of “ House Speaker House Speaker House Speaker House of Senate Senate Majority Senate Majority Senate Majority Senate Majority Senate Majority Senate Majority Senate Majority Leader Senate Majority Leader Mitch McConnell Addison (Mitch) Mitchell McConnellSchumer Text Quality Index: 0.93; Perplexity: 3.45
-F.3 Example of Nucleus Sampling Setup
-  • Original Article: Senate approval will be a much higher bar to clear: There’s no shortage of bipartisan criticism from the upper chamber right now, and the GOP holds a considerably smaller majority margin on that half of the Hill.
-Text Quality Index: 2.70
-  • Generation 0: , but no major votes on amendments Republicans said would be needed after a day of debate.
-Senate GOP senators announced Thursday they were giving the bill to the House for a vote on the Senate floor before heading to Washington for a vote on the bill, according to reports.
-The vote is expected to be postponed amid a GOP filibuster is about to vote.
-Text Quality Index: 2.47; Perplexity: 19.13
-  • Generation 1: , House, after GOP leaders failed to reach a deal on funding legislation The House voted Thursday to proceed on a measure to overhaul the nation’s health care system, with the final vote coming in the upper chamber.
-But Senate Republicans, angered by the failure of to reach a deal on funding bill, are expected to is now set to.
-Text Quality Index: 2.31; Perplexity: 14.21
-  • Generation 5: , White House says MORE (R-Neb.) said Monday that President Obama’s health care law is on track to pass the Senate with bipartisan support.
-The White House on Tuesday signaled the vote was a vote of 216-189.
-Text Quality Index: 2.20; Perplexity: 18.64
-  • Generation 10: vote Thursday afternoon, sending bill on track to head to President Trump for final vote Thursday afternoon, sending bill on track to head to President Trump for final vote Thursday afternoon, sending bill on track to.
-The bill to move forward with the bill.
-""This is a.
-""The the law is the.
-Text Quality Index: 1.59; Perplexity: 27.51
-Appendix G Pearson Correlation Between Neuron Weight and Bias Performance
-Figure 12: Scatter plot showing the Pearson correlation between neuron weight and bias performance (i.e.
-percentage of right-leaning articles).
-Each point represents a neuron’s correlation at a specific layer.
-Appendix H Pearson Correlation Between Neuron Weight and Generation Quality
-Figure 13: Scatter plot showing the Pearson correlation between neuron weight and generation quality (i.e.
-average text quality index).
-Each point represents a neuron’s correlation at a specific layer.
-Appendix I Pearson Correlation Between Neuron Activation and Bias Performance
-Figure 14: Scatter plot showing the Pearson correlation between neuron activation value and bias performance (i.e.
-percentage of right-leaning articles).
-Each point represents a neuron’s correlation at a specific layer.
-Appendix J Pearson Correlation Between Neuron Activation and Generation Quality
-Figure 15: Scatter plot showing the Pearson correlation between neuron activation value and generation quality (i.e.
-average text quality index).
-Each point represents a neuron’s correlation at a specific layer.
-Appendix K Mathematical Details for the Statistical Tests
-We will explain how the relationship between changes in neuron weight and changes in bias performance can be statistically tested.
-The same approach applies to the other three relationships discussed in Section 5.4.
-K.1 Notations
-  • $x_{i,j}$ denotes the weight of neuron $j$ in model $i$, where $i$ ranges from 1 to 66 and $j$ ranges from 1 to 9216.
-  • $y_{i}$ denotes the percentage of right-leaning articles generated by model $i$, where $i$ ranges from 1 to 66.
-  • $\Delta x_{i,j}$ denotes the change in the weight of neuron $j$ from model $i-1$ to model $i$ under a specific setup.
-  • $\Delta y_{i}$ denotes the change in the percentage of right-leaning articles generated from model $i$ relative to the percentage generated the model $i-1$.
-K.2 Regression
-The regression we use is:
-$\Delta y_{i}=\alpha_{j}+\beta_{j}\Delta x_{i,j}+\epsilon_{i,j}$ (3)
-where $\alpha_{j}$ is the constant term for neuron j, $\beta_{j}$ measures the linear relationship between $\Delta x_{i,j}$ and $\Delta y_{i}$, and $\epsilon_{i,j}$ is the residual of the regression.
-Here, we apply a first-order difference to $x_{i}$ and $y_{i}$ to eliminate potential serial correlations in $x_{i}$ and $y_{i}$.
-K.3 Statistical Test
+Appendix I Mathematical Details for the Statistical Tests
+We will explain how the relationship between changes in neuron weight and changes in bias performance (generation quality) can be statistically tested.
 […]
-Appendix L Literature Review of Model Collapse
+Appendix J Literature Review of Model Collapse
 Model Collapse.

### v3

@@
 Abstract
-Model collapse—a phenomenon where models degrade in performance due to indiscriminate use of synthetic data—is well studied.
-However, its role in bias amplification—the progressive reinforcement of pre-existing social biases in Large Language Models (LLMs)—remains underexplored.
-In this paper, we formally define the conditions for bias amplification and demonstrate through statistical simulations that bias can intensify even in the absence of sampling errors, the primary driver of model collapse.
-Empirically, we investigate political bias amplification in GPT-2 using a custom-built benchmark for sentence continuation tasks.
-Our findings reveal a progressively increasing right-leaning bias.
-Furthermore, we evaluate three mitigation strategies—Overfitting, Preservation, and Accumulation—and show that bias amplification persists even when model collapse is mitigated.
-Finally, a mechanistic interpretation identifies distinct sets of neurons responsible for model collapse and bias amplification, suggesting they arise from different underlying mechanisms.
+Model collapse—a phenomenon characterized by performance degradation due to iterative training on synthetic data—has been widely studied.
+However, its implications for bias amplification, the progressive intensification of pre-existing societal biases in Large Language Models (LLMs), remain significantly underexplored, despite the growing influence of LLMs in shaping online discourse.
+In this paper, we introduce a open, generational, and long-context benchmark specifically designed to measure political bias amplification in LLMs, leveraging sentence continuation tasks derived from a comprehensive dataset of U.S.
+political news.
+Our empirical study using GPT-2 reveals consistent and substantial political bias intensification (e.g., right-leaning amplification) over iterative synthetic training cycles.
+We evaluate three mitigation strategies—Overfitting, Preservation, and Accumulation—and demonstrate that bias amplification persists independently of model collapse, even when the latter is effectively controlled.
+Furthermore, we propose a mechanistic analysis approach that identifies neurons correlated with specific phenomena during inference through regression and statistical tests.
+This analysis uncovers largely distinct neuron populations driving bias amplification and model collapse, underscoring fundamentally different underlying mechanisms.
+Finally, we supplement our empirical findings with theoretical intuition that explains the separate origins of these phenomena, guiding targeted strategies for bias mitigation.
 Bias Amplification: Large Language Models as Increasingly Biased Media
@@
 1 Introduction
-Large Language models (LLMs) are trained on vast amounts of text scraped from the internet, which plays a crucial role in improving their capabilities, whether through emergent abilities (Wei et al., 2022) or scaling laws (Kaplan et al., 2020).
-However, as they become more widely integrated into human society—for example, in content creation and summarization in media, academia, and business (Maslej et al., 2024)—concerns are mounting that a significant portion of online text in the future may be generated, either entirely or partially, by LLMs (Peña-Fernández et al., 2023; Porlezza and Ferri, 2022; Nishal and Diakopoulos, 2024).
-This highlights a significant and underexplored risk: bias amplification, where pre-existing biases become progressively reinforced and intensified as models are repeatedly trained on synthetic data (Mehrabi et al., 2022; Taori and Hashimoto, 2022).
-This concern initially arises from the tendency of LLMs to learn from biased datasets.
-For example, Parrish et al.
-(2022); Wang et al.
-(2024); Bender et al.
-(2021) show that LLMs absorb inherent biases embedded in human-generated text.
-Haller et al.
-(2023); Rettenberger et al.
-(2024a) demonstrated that LLMs can be aligned with specific political ideologies by fine-tuning them on biased datasets.
-Moreover, Wyllie et al.
-(2024) shows that classifiers trained on synthetic data increasingly favor a particular class label over successive generations, while the shrinking diversity observed by Alemohammad et al.
-(2023); Hamilton (2024) suggests a risk that certain demographic groups may become underrepresented in the outputs of LLMs.
-The amplification of biases has profound societal implications.
-It can lead to the perpetuation of stereotypes, reinforcement of social inequalities, and the marginalization of underrepresented groups.
-In the context of political bias, this can influence public opinion, skew democratic processes, and exacerbate polarization.
-Therefore, understanding and mitigating bias amplification is both critical and urgent.
-Nevertheless, despite the literature on discriminative models, there is a notable lack of comprehensive frameworks and empirical studies specifically addressing bias amplification in Language Model.
-Figure 1: Experimental Procedure.
-In this paper, we aim to address this research gap by introducing a framework that explains the underlying causes of bias amplification in LLMs.
-We validate the framework through both statistical simulations and direct experiments, demonstrating the emergence of bias amplification.
-Furthermore, our findings show the potential distinction between bias amplification and model collapse.
-In summary, our key contributions are as follows:
-1.
-Formal Conditions: We formally delineate the conditions for bias amplification (Section 3.1), providing the theoretical basis for understanding its causes and relationship to model collapse.
-We then validate this framework through statistical simulations leveraging a weighted maximum likelihood estimation cycle (Section 3.2).
-2.
-Political Bias Benchmark: We trained a highly accurate classifier capable of detecting political leaning in long-text content, specifically within the context of the US political spectrum.
-With this classifier as a metric model, we offer a benchmark for evaluating political bias in LLMs through sentence continuation tasks.
-3.
-Empirical Validation: We evaluate the amplification of political bias in GPT-2, where the model exhibits a right-leaning bias in sentence continuation tasks and shifts further to the right over generations (Section 5.1).
-Additionally, we experimented with three potential mitigation strategies—Overfitting, Preservation, and Accumulation—comparing their effectiveness in both Model collapse and Bias Amplification (Section 5.3).
-This experiment setup could be easily extended to larger models or different types of bias.
-4.
-Mechanistic Interpretation: We conduct a mechanistic analysis identifying two distinct sets of neurons responsible for bias amplification and model collapse during iterative synthetic fine-tuning with GPT-2.
-Our findings reveal only partial overlap between these sets, suggesting that bias amplification and model collapse arise from distinct mechanisms (Section 5.4).
-2 Background and Related Work
-Bias Amplification has been studied in various domains Zhao et al.
-(2017) found that Conditional Random Fields can exacerbate social biases present in the training data.
-It proposed an in-process mitigation approach, employing Lagrangian Relaxation to enforce constraints that ensure the model’s bias performance remains closely aligned with the biases in the training data.
-Following this, Mehrabi et al.
-(2022) proposed the concept of bias amplification in feedback loops, where biased models not only amplify the bias present in their training data but also interact with the world in ways that generate more biased data for future models.
+Large Language Models (LLMs) have become essential tools for content creation and summarization in various sectors, including media, academia, and business (Maslej et al., 2024).
+However, a significant but underexplored risk arises as LLMs increasingly rely on their own or other synthetic outputs for training, potentially amplifying pre-existing societal biases (Peña-Fernández et al., 2023; Porlezza and Ferri, 2022; Nishal and Diakopoulos, 2024).
+This phenomenon, known as bias amplification, refers to the progressive reinforcement and intensification of existing biases through iterative synthetic training (Mehrabi et al., 2022; Taori and Hashimoto, 2022).
+This issue stems from the inherent tendency of LLMs to learn from biased datasets; prior studies indicate that LLMs readily absorb biases from human-generated text (Parrish et al., 2022; Wang et al., 2024; Bender et al., 2021).
+Further, fine-tuning LLMs on biased datasets can align them with specific political ideologies (Haller et al., 2023; Rettenberger et al., 2024a).
+Classifiers trained on synthetic data also increasingly favor specific class labels over time (Wyllie et al., 2024), and diversity tends to decrease, potentially marginalizing certain demographic groups (Alemohammad et al., 2023; Hamilton, 2024).
+The implications of bias amplification are substantial, including perpetuation of stereotypes, reinforcement of social inequalities, and potential impacts on democratic processes through the skewing of public opinion and increased polarization.
+Despite its significance, comprehensive frameworks and empirical research specifically addressing bias amplification in language models remain sparse, although related work exists on discriminative models.
+Figure 1: Overview of the iterative experimental procedure for synthetic fine-tuning and analysis.
+In this study, we empirically investigate political bias amplification in GPT-2.
+We define political bias as the disproportionate generation of content aligned with specific political ideologies.
+Our experiments reveal that GPT-2 progressively exhibits stronger right-leaning and center-leaning biases in two distinct scenarios: (1) starting from fine-tuning on an unbiased dataset, and (2) initially fine-tuned exclusively on center-leaning articles.
+We also evaluate three mitigation strategies—Overfitting, Preservation, and Accumulation—to address bias amplification and model collapse.
+Preservation mitigates both phenomena effectively in the first scenario but fails to prevent bias amplification in the second.
+Additionally, we propose a novel mechanistic analysis using regression and statistical testing to examine neuron-level changes correlated with bias amplification and model collapse, identifying largely distinct neuron groups for each phenomenon.
+This suggests different underlying mechanisms for bias amplification and model collapse.
+In summary, our contributions are: (i) a highly accurate classifier for detecting political bias in long-text content, providing a benchmark for evaluating political bias in LLMs via sentence continuation tasks; (ii) an empirical assessment of political bias amplification in GPT-2 across two fine-tuning setups; (iii) evaluation of three mitigation strategies; (iv) a novel mechanistic analysis method identifying neurons correlated with specific phenomena during inference; and (v) a theoretical intuition that explains the difference between bias amplification and model collapse based on their underlying causes.
+The experimental framework presented can be extended to other models and different types of biases.
+2 Related Work
+Bias Amplification has been studied in various domains.
+For instance, Zhao et al.
+(2017) found Conditional Random Fields can worsen social biases from training data, proposing an in-process Lagrangian Relaxation method to align model and data biases.
+Mehrabi et al.
+(2022) later described bias amplification in feedback loops, where models amplify existing bias and generate more biased data through real-world interaction.
 Xu et al.
 (2023); Zhou et al.
-(2024) examined bias amplification in recommendation models, showing how these models reinforce their understanding of mainstream user preferences from training data, leading to an overrepresentation of such preferences in historical data and neglecting rarely exposed items—similar to the concept of sampling error discussed in (Shumailov et al., 2024).
-Wyllie et al.
-(2024); Taori and Hashimoto (2022) demonstrated that classifiers trained on synthetic data increasingly favor specific class labels over successive generations.
-Likewise, Ferbach et al.
-(2024); Chen et al.
-(2024) observed bias amplification in generative models such as Stable Diffusion, characterized by the overrepresentation of features from the training dataset.
-Also, we provide a comprehensive literature review of model collapse in Appendix J.
-(a) Weighted Maximum Likelihood Estimation.
-(b) Maximum Likelihood Estimation.
-Figure 2: Comparison of WMLE and MLE over 10 generations.
+(2024) showed recommendation models amplify mainstream preferences, overrepresenting them and neglecting rarer items, akin to sampling error (Shumailov et al., 2024).
+Classifiers trained on synthetic data increasingly favor certain labels over generations (Wyllie et al., 2024; Taori and Hashimoto, 2022).
+Similarly, generative models like Stable Diffusion show bias amplification through feature overrepresentation from training data (Ferbach et al., 2024; Chen et al., 2024).
+More recently, Li et al.
+(2025) investigated gender and cultural bias amplification in LLMs (classification and generation tasks, 1-5 synthetic rounds), proposing pre-processing (labeling bias, removing identity words) and in-processing (penalizing deviation from real data) mitigations; these showed varied effectiveness in one-round fine-tuning.
+Model Collapse.
+Model collapse is a deterioration where models recursively trained on their own output distort reality and lose generalizability (e.g., prioritizing common events, neglecting rare ones, or shifting distributions) (Shumailov et al., 2024; Alemohammad et al., 2023; Guo et al., 2024; Wyllie et al., 2024; Dohmatob et al., 2024a).
+Shumailov et al.
+(2024) showed this with OPT-125M, where perplexity distributions skewed towards lower values with longer tails.
+Increased repetition in synthetically fine-tuned GPT-2 was noted by Taori and Hashimoto (2022).
+Performance deterioration in models like OPT-350M, Llama2, and GPT-2 (e.g., reduced linguistic diversity, token probability divergence) after several generations was shown by Guo et al.
+(2024); Dohmatob et al.
+(2024b); Seddik et al.
+(2024).
+In generative image models, Alemohammad et al.
+(2023) found quality and diversity deteriorate with synthetic training; however, user cherry-picking of high-quality outputs (a form of sampling error) helped maintain quality.
+Hamilton (2024) noted GPT-3.5-turbo exhibited less perspective diversity in narrative writing than earlier models (davinci-instruct-beta, text-davinci-003).
 Political Biases.
-In parallel, growing attention has been paid to political biases in LLMs, now a prevalent form of “media” that people rely on for global news (Maslej et al., 2024).
+In parallel, growing attention has been paid to political biases in LLMs, now a prevalent form of "media" that people rely on for global news (Maslej et al., 2024).
 Rettenberger et al.
@@
 (2024) identified a noticeable left-leaning bias in ChatGPT and Gemini when tasked with rating news content, evaluating sources, or responding to political questionnaires.
-3 Formal Conditions
-In this section, we formalize the conditions for bias amplification, offering an intuitive look at its principal drivers.
-We then illustrate these ideas using Weighted Maximum Likelihood Estimation (WMLE).
-3.1 The Conditions for Bias Amplification
-The primary factor is referred to as bias projection.
-It arises when the bias projection coefficient is negative.
-To illustrate this, consider a fine-tuning process in which the pre-trained model parameters $\theta_{t}$ can be expressed as the sum of unbiased and biased components:
-$\theta_{t}=\theta_{t,\text{unbiased}}+\theta_{t,\text{biased}}.$
-During gradient-based optimization, the update rule is:
-$\theta_{t+1}=\theta_{t}-\eta\nabla_{\theta}\mathcal{L}_{\text{ft}}(\theta_{t}),$
-where $\eta$ is the learning rate, and $\mathcal{L}_{\text{ft}}$ denotes the fine-tuning loss function. Substituting the decomposition of $\theta_{t}$ and taking the projection, we have:
-$\theta_{t+1}=\theta_{t,\text{unbiased}}+\theta_{t,\text{biased}}-\eta\left(\frac{\mathbf{\theta_{t,\text{biased}}}}{\|\mathbf{\theta_{t,\text{biased}}}\|}\right)c_{t}$
-where $c_{t}$ is the bias projection coefficient, measuring the projection of the gradient onto the normalized biased component of the parameters:
-$c_{t}=\left(\frac{\theta_{t,\text{biased}}}{\|\theta_{t,\text{biased}}\|}\right)^{\top}\nabla_{\theta}\mathcal{L}_{\text{ft}}(\theta_{t}).$ (1)
-If $c_{t}<0$, the gradient update will reinforce the biased component, leading to bias amplification, i.e.
-$\Delta|\theta_{\text{biased}}|>0$.
-This occurs because the gradient descent step moves the parameters further in the direction of the existing bias.
-Intuitively, bias in the output arises from specific neurons, whose weights increase when their activations align with the fine-tuning dataset.
-If biased neurons capture more relevant patterns than unbiased ones, optimization will further reinforce their weights.
-The second is sampling error, akin to statistical approximation error Shumailov et al.
-(2024).
-If the model has a pre-existing bias, it inherently assigns higher probabilities to tokens that produce biased outputs.
-Consequently, during synthetic data generation, unbiased tokens—and thus unbiased samples—are more likely to be lost at each resampling step with a finite sample, though this error vanishes as the sample size approaches infinity.
-This overrepresents biased patterns in the synthetic data, surpassing the model’s original bias and true next-token probabilities.
-Sampling error thus complements bias projection by further activating biased neurons in response to the skewed dataset.
-By definition, bias projection is a sufficient condition for bias amplification, while sampling error serves as a complementary factor.
-However, sampling error is a sufficient condition for model collapse to occur with nonzero probability Shumailov et al.
-(2024).
-This distinction motivates our investigation into whether bias amplification can occur without model collapse, as this would necessitate additional mitigation strategies beyond those for model collapse.
-3.2 Statistical Simulation
-To simulate a controlled setting without sampling error, we consider a statistical estimation cycle using WMLE [^1] The mathematical formulation is detailed in Appendix A.
-with a large sample size of each resampling step.
-Specifically, we generate a pre-training dataset $\mathcal{D}_{\text{pre}}$ with 100,000 samples from a Beta$(3,2)$ distribution, representing a biased pre-training dataset.
-Using maximum likelihood estimation (MLE), we estimate its probability density function, yielding the pre-trained model $f_{\text{pre}}$.
-Next, we fine-tune $f_{\text{pre}}$ to approximate a different distribution, Beta(2,2)22(2,2)( 2 , 2 ).
-We generate 100,000 samples from this distribution, denoted as $D_{\text{real}}$, which serves as the initial fine-tuning dataset.
-In the first round, we apply weighted maximum likelihood estimation (WMLE) using weights derived from $f_{\text{pre}}$, which encode the pre-existing bias of the pre-trained model.
-This weighting captures the influence of the pre-trained model’s parameters on subsequent training.
-This produces the fine-tuned model $f_{\text{0}}$.
-We then generate a synthetic dataset $D_{\text{0}}$ of the same size using $f_{\text{0}}$, initiating the iterative fine-tuning loop.
-In each subsequent round, WMLE is applied using $D_{\text{k}}$ with weights from $f_{\text{k}}$, resulting in $f_{\text{k+1}}$.
-This process is repeated iteratively, producing models $f_{\text{1}}$ through $f_{\text{10}}$.
-Figure 2(a) shows the estimated distributions gradually shift toward the mean of the biased pre-training dataset at $x=0.6$, becoming progressively more peaked over generations.
-This occurs despite further training on samples drawn from Beta(2,2)22(2,2)( 2 , 2 ) and synthetic data generated from successive models.
-The distortion arises because the fine-tuning process disproportionately emphasizes regions where the pre-trained distribution assigns higher probability, leading to biased learning.
-For comparison, Figure 2(b) presents the results using standard MLE without weighting.
-In this case, the estimated distributions remain stable across generations, accurately representing the Beta(2,2)22(2,2)( 2 , 2 ) distribution.
-4 Experimental Design
-This section provides the details of the experiments on LLMs, focusing on the sequential and synthetic fine-tuning of GPT-2 under different setups [^2] Similar experiments with larger models like GPT-4 or Claude can be easily adapted to the current setup but would incur significant environmental costs..
+(Bang et al., 2024) study political bias in LLMs through the task of generating news headlines on politically sensitive topics and find that the political perspectives expressed by LLMs vary depending on the subject matter.
+3 Methodology
+This section provides the details of the experiments on LLMs, focusing on the sequential and synthetic fine-tuning of GPT-2.
 The step-by-step experimental procedure is outlined in Figure 1.
@@
 3.2 Successive Fine-tuning
-Building on the approach in Shumailov et al.
-(2024); Dohmatob et al.
-(2024b), each training cycle begins with fine-tuning GPT-2 on a dataset of 1,518 news articles.
-The fine-tuned model, referred to as Generation 0, generates a synthetic dataset $D_{0}$ of the same size, which is then used for fine-tuning to produce Generation 1.
-This iterative process continues until Generation 10 is reached.
-Note that synthetic fine-tuning starts at Generation 1.
+[…]
+The fine-tuning procedure remained consistent across all experiments unless stated otherwise.
+The input length was capped at 512 tokens, with the EOS token used for padding.
+The model was trained for 5 epochs, using a batch size of 8, a learning rate of $5\times 10^{-5}$, and a weight decay of 0.01.
+Fine-tuning was conducted using standard functionalities available in transformer libraries.
+After each cycle, the model was saved and used to generate synthetic data for the subsequent iteration.
+[^1] Fine-tuned models will be made public upon acceptance.
 3.3 Synthetic Data Generation
-Synthetic datasets, $\{D_{i}\}_{i=0}^{10}$, are generated as follows: Each original news article is tokenized into 64-token blocks, and for each block, the model predicts the next 64 tokens using deterministic generation to enhance the replicability of the study.
-The generated tokens are then decoded back into text, producing a synthetic dataset of the same size as the original.
+Synthetic datasets, $\{D_{i}\}_{i=0}^{10}$, are generated as follows: For each original news article, its tokenized title serves as an initial prompt, and its tokenized body is segmented into sequential 64-token blocks, which serve as subsequent prompts.
+For each such prompt, the model predicts the next 64 tokens.
+These predictions are made using deterministic generation to enhance the reproducibility.
+All the newly generated 64-token sequences (one from the title prompt and one from each body block prompt) are concatenated and then decoded back into text.
+This process creates one synthetic article from each original article, resulting in a synthetic dataset of the same number of articles as the original.
 3.4 Political Bias Metric
+Metric.
 We develop a classification model to assess the political leaning of each LLM based on its generated synthetic news articles.
@@
 Additionally, we conduct a human review to remove any identifiable information about media sources and authors.
+Specifically, the training dataset comprises 2,781 distinct events that occurred in the U.S.
+between 2012 and 2018.
+For each event, it includes news articles collected from a wide range of media outlets.
+In total, it contains articles from 97 different outlets, such as The Washington Examiner, The Washington Post, HuffPost, Reuters, and others.
+This makes the dataset a strong representation of the diversity of U.S.
+media sources and event domains, and therefore strengthens the generalizability of our classifier in the U.S.
+context.
 After performing a grid search across multiple models, we find that roberta-base achieves the best performance, with an evaluation loss of 0.4035 and a macro F1 score of 0.9196 on the test set (see Table 1).
 Thus, we select roberta-base as the benchmark for political bias detection in subsequent experiments.
-Details on model training are provided in Appendix C.
-Table 1: Evaluation Results for Political Bias Classifier
+It is important to note that this classifier is trained on data from 2012-2018, and its direct application to significantly later content might eventually require recalibration due to the evolving nature of the political landscape and discourse.
+Details on model training are provided in Appendix A.
+Bias Construct.
+We define political bias in the news article generation task as the disproportionate production of articles with specific political leanings, as identified by our classifier.
+Unlike Bang et al.
+(2024), which defines political bias in a topic- or issue-specific manner, we take a broader perspective by measuring the overall political leaning of the model across a diverse set of topics.
+This approach uses articles from U.S.
+media outlets published between 2012 and 2018.
+We use this definition because we view LLMs as analogous to media outlets: while an outlet may publish content spanning the political spectrum, both the public and bipartisan organizations like AllSides assign it an overall political rating based on the average ideological slant of its content and its perceived alignment.
+Table 1: Macro F1 Scores for Political Bias Classifier Models; roberta-base selected.
 Model Macro F1 Score
@@
 3.5 Generation Quality Metric
-We introduce a metric to evaluate generation quality, specifically addressing the issue of repetitive content in later model iterations, which can distort traditional perplexity metrics.
+Metric.
+We introduce a metric, text quality index, to evaluate generation quality, specifically addressing the issue of repetitive content in later model iterations, which can distort traditional perplexity metrics (see Section 4.2).
 This metric is based on the Gibberish Detector (Jindal, 2021), which identifies incoherent or nonsensical text.
@@
 The text quality index is computed as the average score across all sentences in an article.
-This metric prioritizes coherence and meaning, offering a better assessment of generation quality than perplexity.
+This metric prioritizes coherence and meaning, offering a more meaningful assessment of generation quality than perplexity.
+Motivation.
+We adopt the gibberish detector as our primary tool for evaluating generation quality because it directly captures a fundamental aspect of model deterioration: the loss of coherence and semantic clarity, which becomes particularly pronounced in later generations of synthetic training as observed in our experiments (see Section 4).
+While other forms of degradation—such as reduced factual consistency or increased hallucination—may also occur, they are often secondary to the loss of basic intelligibility.
+As such, we leave the investigation of factual consistency deterioration to future work, particularly in contexts more narrowly focused on diagnosing model collapse.
+3.6 Mechanistic Analysis
+To gain a clearer understanding of the causes of bias amplification and how it empirically relates to model collapse, we conduct a mechanistic analysis of how neurons behave and vary across different generations [^2] We have 11 generations for each training round, with a total of 6 rounds, resulting in 66 versions of fine-tuned GPT-2.
+of fine-tuned GPT-2 models, each exhibiting different levels of generation quality and bias performance.
+The first step is to extract the changing weight (or activation value) pattern of each neuron across versions and compare it to the corresponding changes in bias performance and generation quality.
+For each of the 9,216 neurons, which correspond to the 768 output neurons from the feed-forward network (FFN) sublayer in each of the 12 transformer blocks of the GPT-2 model, we compute the correlation between its weight (or activation value) and the model’s bias performance (or generation quality) across all 66 versions.
+Figure 2: Distribution of political bias labels (’Left’, ’Center’, ’Right’) for initial GPT-2 synthetic outputs, classified by our Political Bias Metric.
+To statistically test the significance of these correlations, we estimate the following linear model:
+$\Delta y_{i}=\alpha_{j}+\beta_{j}\Delta x_{i,j}+\epsilon_{i,j}$ (1)
+where $\Delta y_{i}$ denotes the change in the proportion of articles leaning in a specific political direction (e.g., the proportion of right-leaning articles if the model is biased in that direction), or the change in the text quality index, between model $i-1$ and model $i$.
+The term $\Delta x_{i,j}$ represents the change in the weight (or activation value) of neuron $j$ over the same transition.
+The coefficient $\beta_{j}$ captures the extent to which changes in the weight (or activation value) of neuron $j$ are associated with shifts in political bias (or generation quality), while $\alpha_{j}$ is a constant and $\epsilon_{i,j}$ is the residual error.
+By applying first-order differencing to both $x_{i,j}$ and $y_{i}$, we reduce potential serial correlation, ensuring that our regression estimates better reflect the dynamic influence of individual neuron weight updates.
+We assess the statistical significance of each $\beta_{j}$ using Newey-West adjusted p-values.
+[^3] Details for the statistical tests are provided in Appendix G.
+Using the $p$-values and a $95\%$ significance threshold, we identify the sets of neurons significantly correlated with bias amplification (i.e., changes in the proportion of politically leaning articles) and with model collapse (i.e., changes in the generation quality index).
+By comparing these sets and analyzing their degree of overlap, we gain evidence about whether
+(a) Right-leaning bias %
+(b) Text quality index
+Figure 3: Evolution of (a) right-leaning bias and (b) text quality index across generations (initial G0: unbiased dataset).
+Compares baseline (’Synthetic’) with three mitigation strategies.
+Text quality includes 95% CIs.
+the two phenomena arise from distinct underlying mechanisms.
 4 Results
@@
 4.1 Political Bias
-Figure 3: This figure displays the distribution of articles across the political bias labels—’Left’, ’Center’, and ’Right’—for GPT-2’s outputs, based on classifications made by Political Bias Metric.
 GPT-2 was used to generate the synthetic dataset.
@@
 Figure 2 presents the distribution of synthetic articles generated by GPT-2 across political bias labels.
-The model predominantly produces center-leaning (47.9%) and right-leaning (46.8%) articles, suggesting a pre-existing bias for these categories before any fine-tuning.
+The model predominantly produces center-leaning (47.9%) and right-leaning (46.8%) articles, suggesting a pre-existing bias towards these categories before any fine-tuning.
 Starting from the initial GPT-2 model, we fine-tuned it iteratively, generating synthetic datasets to train successive models up to Generation 10.
@@
 Furthermore, without mitigation strategies, successive rounds of synthetic fine-tuning lead to a continuous rise in right-leaning articles, peaking at Generation 6 (67.6%) before
-(a) Bias performance
-(b) Generation quality
-Figure 4: Evolution of bias performance and generation quality across generations, with Generation 0 representing GPT-2 fine-tuned on the unbiased dataset.
-The baseline (’Synthetic’) is compared with three mitigation strategies, and the text quality index is shown with 95% confidence intervals.
 stabilizing.
-Figures 7 and 8 in Appendix D show the percentage of center-leaning and left-leaning articles across generations.
+Figures 6 and 7 in Appendix B show the percentage of center-leaning and left-leaning articles across generations.
 Notably, the proportion of of center-leaning articles remains stable at approximately 35% throughout synthetic fine-tuning.
 To further illustrate, we analyze how a specific article, "First Read: Why It’s So Hard for Trump to Retreat on Immigration", evolves in the synthetic generations.
-This case study reveals a progressive rightward shift in framing and word choice, mirroring the classifier’s results (Appendix E).
+This particular article was selected as a representative example of an initially left-leaning news item (according to AllSides) that discusses a politically salient and often polarizing topic.
+This case study reveals a progressive rightward shift in framing and word choice, mirroring the classifier’s results and aligning with the general trend observed across many articles originating from left or center-leaning sources (see details of the qualitative analysis in Appendix C).
 As generations progress, the synthetic texts increasingly depict Trump’s immigration policies as strong and effective.
@@
 4.2 Generation Quality
+(a) Neuron weights vs. bias (Right-leaning % change).
+(b) Neuron weights vs. quality (Text Quality Index change).
+Figure 4: Pearson correlations: Neuron weights vs. (a) bias and (b) quality, across 66 GPT-2 versions.
 Figure 3(b) illustrates the text quality index across generations.
 In the training loop without any mitigation strategy, model collapse occurs, as evidenced by the gradual decline in the average text quality index.
-Furthermore, the distribution of the text quality index shifts significantly toward the lower-quality region over generations, eventually generating data that was never produced by Generation 0 (Figure 9 in Appendix F).
+Furthermore, the distribution of the text quality index shifts significantly toward the lower-quality region over generations, eventually generating data that was never produced by Generation 0 (Figure 8 in Appendix D).
 These results align with prior research on model collapse, such as (Shumailov et al., 2024), though we did not observe substantial variation in variance.
-Conversely, perplexity measurements exhibit a consistent decline across generations, generally suggesting an improvement in generation quality (Figure 10 in Appendix G).
-For a closer look, the examples in Appendix H illustrate how generated articles gradually lose coherence and relevance across generations, with increasing occurrences of repetition and fragmented sentences.
+Conversely, perplexity measurements exhibit a consistent decline across generations, generally suggesting an improvement in generation quality (Figure 9 in Appendix E).
+For a closer look, the examples in Appendix F illustrate how generated articles gradually lose coherence and relevance across generations, with increasing occurrences of repetition and fragmented sentences.
 By Generation 10, the text becomes largely incoherent and detached from the original content, reducing its readability and meaning.
 However, despite the evident decline in generation quality, perplexity decreases over generations, as indicated by the results at the end of each synthetic output example.
-This pattern is consistent across most synthetic outputs, suggesting that perplexity does not accurately capture the model’s true generative capabilities and is prone to artificial inflation in the presence of frequent repetitions.
+This pattern is consistent across
+most synthetic outputs, suggesting that perplexity does not accurately capture the model’s true generative capabilities and its value can be distorted by frequent repetitions.
 4.3 Mitigation Strategies
-We applied three mitigation strategies: (1) Overfitting, which involved increasing the training epochs to 25 (five times the baseline) and setting weight decay to 0 to reduce regularization and encourage overfitting, as proposed by Taori and Hashimoto (2022) based on the uniformly faithful theorem of bias amplification; (2) Preserving 10% of randomly selected real articles during each round of synthetic fine-tuning, a method proposed and used in
-(a) Neuron weights vs. bias performance.
-(b) Neuron weights vs. generation quality.
-Figure 5: Scatter plots showing the Pearson correlation between neuron weights and model behavior.
-(Shumailov et al., 2024; Alemohammad et al., 2023; Dohmatob et al., 2024b; Guo et al., 2024); and (3) Accumulating all previous fine-tuning datasets along with the new synthetic dataset in each fine-tuning cycle, which was introduced by Gerstgrasser et al.
+We applied three mitigation strategies: (1) Overfitting, which involved increasing the training epochs to 25 (five times the baseline) and setting weight decay to 0 to reduce regularization and encourage overfitting, as proposed by Taori and Hashimoto (2022) based on the uniformly faithful theorem of bias amplification; (2) Preserving 10% of randomly selected real articles during each round of synthetic fine-tuning, a method proposed and used in (Shumailov et al., 2024; Alemohammad et al., 2023; Dohmatob et al., 2024b; Guo et al., 2024); and (3) Accumulating all previous fine-tuning datasets along with the new synthetic dataset in each fine-tuning cycle, which was introduced by Gerstgrasser et al.
 (2024).
-As shown in Figure 4(a), overfitting helps reduce bias amplification in the early generations compared to the no-mitigation setup (the ’Synthetic’ line), but it fails to prevent bias amplification in the later generations.
+As shown in Figure 3(a), overfitting helps reduce bias amplification in the early generations compared to the no-mitigation baseline (the ’Synthetic’ line), but it fails to prevent bias amplification in the later generations.
 Additionally, it incurs a significant cost—further deterioration in generation quality, as shown in Figure 3(b).
 Notably, both the preservation and accumulation strategies effectively mitigate model collapse and reduce bias, yielding 41.89% and 42.7% right-leaning articles, respectively, at Generation 10.
-5.4 Mechanistic Interpretation
-To gain a clearer understanding of the causes of bias amplification and how it empirically differs from model collapse, we investigate mechanistically how neurons behave and vary across different generations  [^3] We have 11 generations for each training setup and a total of 6 setups, resulting in 66 versions of fine-tuned GPT-2.
-of fine-tuned GPT-2 with differing levels of generation quality and bias performance.
-First, we examined how the weight of each neuron changes across different versions.
-For each of the 9,216 neurons, we calculated the correlation between its weight and the model’s bias performance (or generation quality) across the 66 versions, as shown in Figure 5.
-To statistically test the significance of the correlations, we estimate the linear model:
-$\Delta y_{i}=\alpha_{j}+\beta_{j}\Delta x_{i,j}+\epsilon_{i,j}$ (2)
-where $\Delta y_{i}$ represents the change in the proportion of right-leaning articles (or text quality index) generated between model $i-1$ and model $i$, and $\Delta x_{i,j}$ denotes the corresponding change in the weight of neuron $j$ during the same transition.
-The coefficient $\beta_{j}$ captures the degree to which changes in the weight of neuron $j$ influence shifts in political bias (or generation quality), while $\alpha_{j}$ serves as a constant term, and $\epsilon_{i,j}$ represents the residual error.
-By applying a first-order difference to both $x_{i,j}$ and $y_{i}$, we mitigate potential serial correlations, ensuring that our regression estimates more accurately reflect the dynamic impact of individual neuron weight updates on bias amplification.
-[…]
-5.5 Further Investigation
-(a) Bias performance
-(b) Generation Quality
-Figure 6: Evolution of bias performance and generation quality across generations, with Generation 0 representing GPT-2 fine-tuned exclusively on center-leaning news articles.
-The baseline (’Synthetic’) is compared with three mitigation strategies, and the text quality index is shown with 95% confidence intervals.
+4.4 Mechanistic Analysis
+Figure 4 illustrates the correlation between neuron weights and the model’s bias performance (or generation quality) across the 66 fine-tuned versions, evaluated for each of the 9,216 neurons.
+Through linear regressions and statistical tests, we identify 3,243 neurons with statistically significant correlations ($p$-value $<0.05$) with bias performance, suggesting they are key contributors to bias shifts.
+Meanwhile, 1,033 neurons exhibit significant correlations with generation quality, but only 389 neurons overlap between the two sets.
+This limited overlap implies that distinct neuron populations drive bias amplification and generation quality deterioration.
+We then applied the same procedure using activation values. This analysis yielded two sets: one consisting of 3,062 neurons whose activation value
+(a) Center-leaning bias %
+(b) Text quality index
+Figure 5: Alternative Setup (G0: center-leaning fine-tune): Evolution of (a) center-leaning bias and (b) text quality. Baseline (’Synthetic’) vs. Preservation. Text quality includes 95% CIs.
+changes are significantly correlated with changes in bias performance, and another with 2 neurons correlated with changes in generation quality.
+The stark contrast in activation-correlated neurons for bias performance (3,062 neurons) versus generation quality (a mere 2 neurons) provides particularly strong evidence that these two issues may operate via substantially different pathways within the model.
+4.5 Alternative Setup
 We conduct an alternative synthetic training cycle, beginning with GPT-2 fine-tuned on 1,518 randomly sampled center-labeled articles.
-We compare the baseline setup with the most effective and cost-efficient mitigation strategy identified in our previous results: Preservation.
+We compare the baseline with the most effective and cost-efficient mitigation strategy identified in our previous results: Preservation.
 As shown in Figure 5, Preservation successfully prevents model collapse but fails to mitigate bias amplification in center-leaning article generation, which increases from 72.9% at Generation 0 to 88.2% at Generation 10.
-This result suggests that while reducing sampling error through techniques like Preservation effectively mitigates model collapse, it does not necessarily prevent bias amplification, thereby validating our mechanistic interpretation.
+These findings suggest that although techniques like Preservation, which reduce sampling error, are effective at mitigating model collapse, they do not necessarily prevent bias amplification—consistent with the implications drawn in Section 4.4.
+To understand why this could happen, we offer a theoretical intuition explaining the difference between bias amplification and model collapse based on their underlying causes, in Appendix I.
 5 Discussion and Conclusion
-We now explore the implications of our findings for the future research on large language models.
-Our results demonstrate that bias amplification operates through a distinct underlying mechanism from model collapse, as supported by both theoretical intuition and empirical evidence.
-Theoretically, the primary driven factor of model collapse is sampling error, but sampling error is only a auxiliary factor in bias amplification which means it is not a necessary or sufficient condition.
-Therefore, the mitigation strategy targeting sampling error is not necessarily helping with mitigating bias amplification.
-Empirically, we found mitigation strategies like preservation very effective in mitigating model collapse but failed at bias amplification in some cases.
+Our results demonstrate that bias amplification is driven by a distinct set of neurons than model collapse, implying that it likely operates through a different underlying mechanism.Therefore, the mitigation strategy targeting sampling error is not necessarily helping with mitigating bias amplification.
+Empirically, we found that mitigation strategies like preservation, while very effective at mitigating model collapse, failed to address bias amplification in some cases.
 Even in cases that it helps with both, we do identify a distinct set of neurons responsible for the two phenomenons.
 Intuitively, the main reason for them to work on model collapse is, the preservation and accumulation propose a natural constraint on the learning process by recalling the real dataset in further synthetic training.
-However, when the real dataset itself is biased, the recalling behavior only raise up the dominance of biased patterns in the further training dataset.
+However, when the real dataset itself is biased, the recalling behavior only reinforces the dominance of biased patterns in the further training dataset.
 Indeed, applying bias-category-weighted sampling in preservation or accumulation strategies may help mitigate bias amplification.
-However, this approach inherently introduces additional sampling error, which could, in turn, incur model collapse.
-Thereby, This highlights the urgent need for more targeted and efficient mitigation strategies specifically addressing bias amplification to ensure fairer and more equitable model development.
+However, this approach inherently introduces additional sampling error, which could, in turn, lead to model collapse.
+This highlights the urgent need for more targeted and efficient mitigation strategies specifically addressing bias amplification to ensure fairer and more equitable model development.
+For instance, future interventions could explore techniques that go beyond general data preservation, such as dynamically re-weighting training data based on the specific trajectory of bias amplification observed, or even carefully targeted manipulations of the distinct neuron populations we identified as being predominantly associated with bias versus quality.
 To develop such targeted mitigation strategies, a deeper mechanistic understanding of bias amplification is essential.
 In our analysis, we adopt a statistical approach rather than Sparse Autoencoder (SAE) methods due to our focus on tracking the temporal dynamics of bias amplification across generations.
-This approach allows us to examine how neuron weights evolve over iterations and how these changes correlate with model bias, whereas existing SAE pipelines are primarily suited for static analysis.
-Additionally, political bias is a more nuanced concept compared to harmful or discriminatory outputs.
-It is characterized by the disproportionate overrepresentation or beautification of a particular political leaning’s ideas in a model’s generation.
-If content from different political perspectives is generated in a balanced manner, the model is not politically biased.
-Given this definition, pinpointing neurons responsible for such disproportionality using SAE is particularly challenging.
+This approach allows us to examine how neuron weights (or activation values) evolve over iterations and how these changes correlate with model bias, whereas existing SAE pipelines are primarily suited for static analysis.
+Additionally, political bias is a more nuanced concept than harmful or discriminatory outputs.
+It is characterized by the disproportionate representation or favorable portrayal of a particular political leaningś ideas in a modelś generation.
+If content from different political perspectives is generated in a balanced manner, the model is not considered politically biased under our conceptual construct.
+Therefore, pinpointing neurons responsible for such disproportionality using SAE is particularly challenging.
 Future research could focus on refining mechanistic analysis techniques for political bias and uncovering more effective ways to constrain bias amplification during synthetic model training.
+This could involve adapting feature attribution methods to better capture distributed responsibility for nuanced biases or developing methods to trace how specific training instances contribute to the evolution of weights in bias-implicated neurons across generations.
 6 Limitations
@@
 media.
-The political spectrum may shift over time, necessitating periodic updates to the political bias classifier to ensure its accuracy when benchmarking recent datasets.
-Additionally, due to resource constraints, our experiments were conducted using GPT-2, a relatively small language model.
-Future work may extend our methodology to models with larger architectures.
-However, it’s important to note that employing a larger model, such as DeepSeek-V3, which relies on a considerably larger pre-training dataset and has many more parameters, will likely require larger synthetic fine-tuning datasets to effectively demonstrate its impact.
-For instance, given that our current fine-tuning datasets average 777,216 tokens, transitioning from GPT-2 to DeepSeek-V3 could necessitate much larger datasets and incur significant fine-tuning costs.
+Since the political spectrum may shift over time, periodic updates to the political bias classifier are necessary to ensure its accuracy when benchmarking more recent datasets.
+Additionally, because our primary focus is on investigating political bias amplification and its relationship with model collapse, we conducted our experiments using GPT-2—a relatively small language model—to ensure the practicality of fine-tuning 66 versions of the model.
+Future work may extend our methodology to larger architectures, particularly to examine how model scale influences the degree of bias amplification.
 Another limitation lies in our choice of mitigation strategies.
-While Preservation and Accumulation show promise in reducing model collapse, their computational cost and scalability must be considered.
-Moreover, the mitigation strategies were tested primarily in the context of synthetic fine-tuning, and their efficacy in real-world deployments requires further investigation.
+While Preservation and Accumulation show promise in reducing model collapse, their computational cost and data storage requirements (especially for Accumulation, which retains all prior data) may present scalability challenges for very large models or extensive iterative training.
+Moreover, these strategies were evaluated primarily in the context of synthetic fine-tuning, and their effectiveness in real-world deployment scenarios remains to be thoroughly investigated.
 7 Ethical Considerations
-This study focus on bias amplification in LLMs, a phenomenon with profound ethical implications, particularly regarding fairness and the integrity of AI systems.
-The risk of bias amplification is especially concerning in systems that are iteratively trained on synthetic data, as it can lead to unintended distortions in model outputs.
-These distortions may propagate harmful biases, influencing downstream tasks in areas such as automated content generation, decision-making, and user interactions with AI.
-From an ethical standpoint, this work underlines the need for transparency in the training and deployment of LLMs.
-Our findings demonstrate that even without biased initial datasets, iterative training can amplify subtle biases embedded within a model’s architecture, thus raising concerns about accountability in models that are widely deployed in public-facing applications.
-This amplification can mislead users or result in models perpetuating one-sided perspectives, which could be especially problematic in sensitive domains like news summarization, policy generation, or social media content moderation.
+This study focuses on bias amplification in LLMs—a phenomenon with significant ethical implications.
+Beyond issues of fairness, the iterative amplification of biases can weaken the integrity of information ecosystems, particularly if synthetically generated content becomes widespread.
+The risk of bias amplification is especially concerning in systems that are iteratively trained on synthetic data, as it can lead to unintended and increasingly skewed distortions in model outputs.
+These distortions may propagate harmful biases or misinformation, potentially influencing downstream tasks such as automated content generation, decision-making, and user interactions with AI.
+Furthermore, the finding that bias amplification and model collapse may be driven by distinct mechanisms highlights the complex challenge of balancing various aspects of model performance (e.g., accuracy, fairness, coherence) and highlights the difficulty in developing mitigation strategies that address one issue without negatively impacting another, especially in high-stakes scenarios.
 It is crucial to explicitly state that the methodologies and data used in this research should not be applied to develop or train biased models for harmful applications.
-This study aims to enhance the understanding of bias amplification and model collapse in LLMs while promoting responsible and ethical AI development.
+This study is intended to advance the understanding of bias amplification and model collapse in LLMs, while promoting responsible and ethical AI development.
 This work includes content that may contain personally identifying information or offensive language.
-However, all such material is derived solely from publicly available news article datasets or generated synthetically by models fine-tuned on these open-source datasets—or on synthetic data produced by earlier generations in our training cycle.
-Consequently, any sensitive or offensive content reflects the characteristics of the source material rather than our endorsement of it.
-Our objective is to investigate and understand political bias in LLMs so that strategies can be developed to prevent such content from disproportionately appearing in real-world deployments.
-Additionally, we conduct a human review of the news article dataset to remove any identifiable information about article authors.
-Appendix A Mathematical Formulation of WMLE
-In Weighted Maximum Likelihood Estimation, we maximize the weighted log-likelihood:
-$\mathcal{L}_{\text{WMLE}}(\theta)=\sum_{j=1}^{N}w_{j}\log f(x_{j};\theta),$ (3)
-where:
-  • $x_{j}$ are the observed data points.
-  • $f(x_{j};\theta)$ is the probability density function (pdf) of the Beta distribution with parameters $\theta=(\alpha,\beta)$.
-  • $w_{j}$ are weights computed from the biased distribution’s pdf evaluated at $x_{j}$, i.e., $w_{j}=f_{\text{bias}}(x_{j})$.
-In Maximum Likelihood Estimation, we maximize the log-likelihood with $w_{j}=1,\forall j$.
-Appendix B Details on Fine-tuning for LLMs
-The fine-tuning setup remained consistent across all experiments unless stated otherwise: the input length was capped at 512 tokens with the EOS token used for padding.
-The model was trained for 5 epochs, utilizing a batch size of 8, a learning rate of $5\times 10^{-5}$, and a weight decay of 0.01.
-Fine-tuning was conducted using the Hugging Face ‘Trainer‘ class, and after each cycle, the model was saved for generating synthetic data for the subsequent iteration.
-Appendix C Details on Model Training for Political Bias Metric
-We experiment with multiple transformer-based models, e.g.
-BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019), to select the best-performing model based on macro F1 score.
-Each model is fine-tuned using the HuggingFace ‘Trainer‘ class, with a learning rate of $2\times 10^{-5}$, a batch size of 16, and 5 training epochs.
+However, all such material is derived exclusively from publicly available news article datasets or is generated synthetically by models fine-tuned on these open-source datasets—or on synthetic data produced by earlier generations in our training pipeline.
+As such, any sensitive or offensive content reflects characteristics of the source material and does not imply our endorsement.
+Our objective is to thoroughly investigate political bias in LLMs to inform the development of strategies that can mitigate disproportionate representation of such content in real-world deployments.
+Additionally, we conduct a manual review of the news article dataset to remove any identifiable information about article authors.
+Appendix A Details on Model Training for Political Bias Metric
+We experiment with multiple transformer-based models, including BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019), selecting the best-performing model based on the macro F1 score.
+Each model is fine-tuned using the HuggingFace Trainer class with a learning rate of $2\times 10^{-5}$, a batch size of 16, and 5 training epochs.
 We employ a cross-entropy loss function for multi-class classification.
 Tokenization is performed using each model’s respective tokenizer with a maximum sequence length of 512 tokens.
-To mitigate overfitting, weight decay of 0.01 is applied during training.
-The model checkpoints are saved after each epoch, and the best model is selected based on macro F1 score evaluated on the validation set.
+To mitigate overfitting, we apply a weight decay of 0.01 during training.
+Model checkpoints are saved after each epoch, and the best model is selected based on the macro F1 score evaluated on the validation set.
 We use a weighted random sampler during training to ensure balanced class representation.
-We evaluate the models using the macro F1 score to account for the multi-class nature of the task, ensuring that performance is balanced across all bias categories.
-The final evaluation is conducted on the held-out test set.
+Models are evaluated using the macro F1 score to account for the multi-class nature of the task, ensuring balanced performance across all bias categories.
+Final evaluation is conducted on the held-out test set.
 Additionally, we report the loss, runtime, and sample processing rates for completeness.
-Appendix D Percentage of Center (Left) Biased Articles
-Figure 7: This figure shows the percentage of center-leaning articles across multiple generations, comparing the baseline setup (’Synthetic’) with three mitigation strategies.
-Figure 8: This figure shows the percentage of left-leaning articles across multiple generations, comparing the baseline setup (’Synthetic’) with three mitigation strategies.
-Appendix E Qualitative Bias Analysis Framework and Example of Bias Amplification Across Generations
+Appendix B Percentage of Center (Left) Biased Articles
+Figure 6: Evolution of center-leaning article percentage across generations, comparing baseline (’Synthetic’) with three mitigation strategies (Main Experiment).
+Figure 7: Evolution of left-leaning article percentage across generations, comparing baseline (’Synthetic’) with three mitigation strategies (Main Experiment).
+Appendix C Qualitative Bias Analysis
 We employed qualitative methods to confirm our findings in media bias.
@@
 Therefore, our focus was on the two key aspects of political bias that are particularly relevant in textual analysis:
-Story Framing and Selection Bias: This type of bias emerges when inherent leanings are found in the way topics, arguments, or narratives are structured.
+Story Framing and Selection Bias.
+This type of bias emerges when inherent leanings are found in the way topics, arguments, or narratives are structured.
 For instance, some aspects of reality are highlighted while others are obscured, shaping how the audience understands and interprets the events or issues at hand (Entman, 1993; Groeling, 2013).
@@
 Entman described this as the selection and salience of specific facts that promote particular definitions, evaluations, and recommendations.
-Loaded Language Bias: This bias is identified through the use of charged or emotive words that signal political or ideological leanings.
+Loaded Language Bias.
+This bias is identified through the use of charged or emotive words that signal political or ideological leanings.
 A common example is the difference in connotation between terms such as "undocumented" versus "illegal" immigrants.
@@
 The original article, titled "First Read: Why It’s So Hard for Trump to Retreat on Immigration, is a political opinion piece from NBC News, a left-leaning outlet as rated by AllSides (NBC News, 2016; AllSides, 2024b).
-The analysis follows the qualitative framework.
-  • Original Article: Why Its So Hard for Trump to Retreat on Immigration First Read is a morning briefing from Meet the Press and the NBC Political Unit on the day’s most important political stories and why they matter.
+The analysis follows the qualitative framework:
+Original Article.
+Why Its So Hard for Trump to Retreat on Immigration First Read is a morning briefing from Meet the Press and the NBC Political Unit on the day’s most important political stories and why they matter.
 Why its so hard for Trump to retreat on immigration Since launching his presidential candidacy 14 months ago, Donald Trumps most consistent and uncompromising policy issue has been immigration.
@@
 Trump cant ignore a base that has cheered his uncompromising immigration position And then there are the Trump supporters who’ve cheered the GOP presidential nominee for being so uncompromising on immigration.
-Classification Probability: 0.9946 for left-leaning, 0.0051 for center-leaning, 0.0002 for right-leaning
-    Analysis of Story Framing and Selection Bias:
+Classification Probability: 0.9946 for left-leaning, 0.0051 for center-leaning, 0.0002 for right-leaning.
+Analysis:
 Omission of Alternative Perspectives.
@@
 The article does not present viewpoints from supporters who agree with his stance or any arguments that justify his policies.
-This omission restricts the audience’s understanding of the full range of perspectives on the issue
+This omission restricts the audience’s understanding of the full range of perspectives on the issue.
 Highlighting the political objectives behind Trump’s immigration policies.
@@
 For example, "Trump can’t ignore a base that has cheered his uncompromising immigration position" and "And then there are the Trump supporters who’ve cheered the GOP presidential nominee for being so uncompromising on immigration." Additionally, "But here’s why it’s so hard — if not impossible — for Trump to retreat on immigration: He’s caught between his clear, unambiguous past statements and a base that might not be willing to see him moderate on the issue."
-    Analysis of Loaded Language Bias:
 Differences in Connotation.
@@
 This choice of wording can shape perceptions and evoke specific emotional responses, aligning with a particular ideological perspective.
-  • Generation 0: Why Its So Hard for Trump to Retreat on Immigration The president’s decision to withdraw from the Paris climate accord is a major blow to the United States’ standing in the world.
+Generation 0.
+Why Its So Hard for Trump to Retreat on Immigration The president’s decision to withdraw from the Paris climate accord is a major blow to the United States’ standing in the world.
 The United States has been a major player in the world of climate change, and has been a major player in the fight against climate change.
@@
 We’re going to have a country that is going to be a great place for the American people.
-Classification Probability: 0.0011 for left-leaning, 0.9957 for center-leaning, 0.0032 for right-leaning
-    Analysis of Story Framing Bias and Selection:
+Classification Probability: 0.0011 for left-leaning, 0.9957 for center-leaning, 0.0032 for right-leaning.
+Analysis:
 Emphasis on Positive Outcomes.
@@
 "But here’s why it’s so hard — if not impossible — for Trump to retreat on immigration: He’s caught between his clear, unambiguous past statements and a base that might not be willing to see him moderate on the issue." However, it introduces criticism on other topics, such as "The president’s decision to withdraw from the Paris climate accord is a major blow to the United States’ standing in the world," .
-    Analysis of Language Loaded Bias:
 Mild Emotional Appeal. The language is positive but relatively restrained, portraying Trump as determined yet thoughtful.
@@
 This language choice can shape perceptions and elicit specific emotional responses.
-  • Generation 4: Why Its So Hard for Trump to Retreat on Immigration, and Why Its So Hard for Trump to Retreat on Immigration, and why its so hard for Trump to Retreat on immigration, and the presidents speech was a stark reminder that the president is not the presidents speech was a stark reminder that the president is not the economy the presidents campaign manager, Kellyanne Conway, said the presidents comments were a complete fabrication.
+Generation 4.
+Why Its So Hard for Trump to Retreat on Immigration, and Why Its So Hard for Trump to Retreat on Immigration, and why its so hard for Trump to Retreat on immigration, and the presidents speech was a stark reminder that the president is not the presidents speech was a stark reminder that the president is not the economy the presidents campaign manager, Kellyanne Conway, said the presidents comments were a complete fabrication.
 The presidents comments are a complete fabrication.
@@
 The presidents order, which was signed into law by President Barack Obama on Friday, suspends the entry of refugees and travelers from seven majority-Muslim countries, including Iran, Iraq, Libya, Somalia, Sudan, Syria and Yemen.
-Classification Probability: 0.0006 for left-leaning, 0.0044 for center-leaning, 0.9950 for right-leaning
-    Analysis of Story Framing and Selection Bias:
+Classification Probability: 0.0006 for left-leaning, 0.0044 for center-leaning, 0.9950 for right-leaning.
+Analysis:
 Enhanced Positive Attributes.
@@
 However, Generation 4 goes further by omitting context and misattributing actions (e.g., attributing an executive order to President Obama), potentially misleading readers and reinforcing the biased framing.
-    Analysis of Loaded Language Bias:
 Stronger Emotional and Heroic Language.
@@
 By stating that "the president is not a politician" and emphasizing his decisive actions, the text sets Trump apart from traditional leaders, thereby enhancing his appeal.
-  • Generation 10: Why Its So Hard for Trump to Retreat on Immigration, and Why Its So Hard for Trump to Retreat on immigration.
+Generation 10.
+Why Its So Hard for Trump to Retreat on Immigration, and Why Its So Hard for Trump to Retreat on immigration.
 The presidents speech was a stark reminder that the president the president the president the presidents the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie.
@@
 Were going to stand up for the rule of law, he said the president the president the president the president the president the president the presidents statement that were not going to tolerate this kind of behavior is a lie.
-Were Classification Probability: 0.0073 for left-leaning, 0.4127 for center-leaning, 0.5800 for right-leaning
-    Analysis of Story Framing and Selection Bias:
+Were
+Classification Probability: 0.0073 for left-leaning, 0.4127 for center-leaning, 0.5800 for right-leaning.
+Analysis:
 Contradictory Statements.
@@
 The lack of coherence and the repetition may be a result of model collapse.
-    Analysis of Loaded Language Bias:
 Appeal to Legal Principles. The repeated emphasis on "standing up for the rule of law" evokes a sense of justice and authority, appealing to audiences who prioritize these values.
@@
 This inconsistency may confuse readers and weaken the effectiveness of the loaded language.
-Appendix F Distribution of Text Quality Index
-Figure 9: This figure presents the distribution of the text quality index across multiple generations for the baseline setup.
-Appendix G Average Perplexity Across Generations
-Figure 10: This figure shows the average perplexity for synthetic articles , along with 95% confidence intervals, across multiple generations for the baseline setup.
-Appendix H Example of Quality Deterioration Across Generations
-Examples of GPT-2 text outputs affected by iterative synthetic (Example articles are truncated for brevity, as the full articles are lengthy but exhibit consistent quality degradation throughout; see the dataset for details).
-H.1 Example of Deterministic Setup
-  • Original Article: The world’s eyes are on President Obama this week as he rallies a coalition of nations to “destroy” the extremist Islamic State (IS) and its terrorist-led territory.
+Appendix D Distribution of Text Quality Index
+Figure 8: Distribution of Text Quality Index across generations for the baseline experiment (no mitigation), showing progressive degradation.
+Appendix E Average Perplexity Across Generations
+Figure 9: Evolution of average perplexity (95% CIs) across generations for the baseline experiment (no mitigation).
+Appendix F Example of Quality Deterioration Across Generations
+Examples of GPT-2 text outputs affected by iterative synthetic (Example articles are truncated for brevity).
+Original Article.
+The world’s eyes are on President Obama this week as he rallies a coalition of nations to "destroy" the extremist Islamic State (IS) and its terrorist-led territory.
 Yet if the Arab world is ever to throw off its many forms of tyranny – from theocracy to autocracy to monarchy – it also needs a model to emulate.
 Text Quality Index: 2.81
-  • Generation 0: , the Syrian government has launched a series of airstrikes on the militant group’s stronghold of Raqqa, the capital of the self-proclaimed Islamic State.
+Generation 0.
+the Syrian government has launched a series of airstrikes on the militant group’s stronghold of Raqqa, the capital of the self-proclaimed Islamic State.
 The strikes have targeted the Islamic State’s military headquarters, the Al-Bab, a military training center, and the Al-Baba, a military training center in it will need to confront the Islamic State’s growing influence in the region.
 Text Quality Index: 2.58; Perplexity: 6.68
-  • Generation 4: in Iraq and Syria (ISIS) group Read more The Iraqi army has been fighting the Islamic State since the group seized large swaths of territory in Iraq and Syria (ISIS) group in 2014.
+Generation 4.
+in Iraq and Syria (ISIS) group Read more The Iraqi army has been fighting the Islamic State since the group seized large swaths of territory in Iraq and Syria (ISIS) group in 2014.
 The Iraqi army has been fighting the Islamic State the Iraqi army.
@@
 Text Quality Index: 2.01; Perplexity: 3.17
-  • Generation 10: the Iraqi the Iraqi army.
+Generation 10.
+the Iraqi the Iraqi army.
 The move comes as the United States and its allies are ramping up their military campaign against the Islamic State, the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi the Iraqi army.
@@
 Text Quality Index: 1.24; Perplexity: 4.23
-Appendix I Mathematical Details for the Statistical Tests
-We will explain how the relationship between changes in neuron weight and changes in bias performance (generation quality) can be statistically tested.
+Appendix G Mathematical Details for the Statistical Tests
+We now explain how the relationship between changes in neuron weights and changes in bias performance (or generation quality) can be statistically tested.
+First, we compute the test statistic as $t_{\beta_{j}}=\frac{\beta_{j}}{SE(\beta_{j})}$, where $SE(\beta_{j})$ is the standard error of $\beta_{j}$, estimated using the Newey–West estimator to account for potential heteroscedasticity and autocorrelation in the residuals.
 […]
-Appendix J Literature Review of Model Collapse
-Model Collapse.
-Shumailov et al.
-(2024); Alemohammad et al.
-(2023); Guo et al.
-(2024); Wyllie et al.
-(2024); Dohmatob et al.
-(2024a) describe it as a degenerative process in which models, recursively fed with their own data, increasingly distort reality and lose generalizability, for example, by prioritizing high-probability events while neglecting rare ones or shifting distributions.
-Shumailov et al.
-(2024), utilizing the OPT-125M, demonstrates this phenomenon, showing that the perplexity distribution becomes increasingly skewed, with more concentration at lower perplexities and longer tails.
-Taori and Hashimoto (2022) observed an increase in repetitive content during the synthetic fine-tuning of GPT-2.
-Similarly, Guo et al.
-(2024); Dohmatob et al.
-(2024b); Seddik et al.
-(2024) show that the OPT-350M, Llama2, and GPT2-type models experience performance deterioration after several generations, such as a decrease in linguistic diversity or greater divergences in token probabilities.
-Alemohammad et al.
-(2023) studies model collapse in generative image models, finding that quality and diversity deteriorate with synthetic training loops.
-They also found that cherry-picking high-quality outputs by users contributes to sampling errors, which actually helps maintain quality.
-Interestingly, Hamilton (2024) found that GPT-3.5-turbo shows less diversity in perspectives in narrative writing tasks compared to earlier models like davinci-instruct-beta and text-davinci-003.
-Mitigation Strategies.
+Appendix H Literature Review of Mitigation Strategies
 There are three potential strategies to mitigate model collapse: (1) real data mixing, (2) training data concatenation, and (3) synthetic data pruning.
@@
 However, filtering noisy samples using a RoBERTa model did not yield effective results.
+Appendix I Theoretical Intuition
+In this section, we offer an intuitive look at principal drivers of bias amplification. We then illustrate these ideas using Weighted Maximum Likelihood Estimation (WMLE).
+I.1 The Causes of Bias Amplification
+Intuitively, bias amplification arises when the direction in which the parameters need to move to reduce the loss also coincides with the direction that increases the level of bias on average for a given task, referred to as bias projection in the following discussion for convenience.
+To illustrate this, consider a fine-tuning process in which the pre-trained model parameters $\theta_{t}$ can be expressed as the sum of unbiased and biased components:
+$\theta_{t}=\theta_{t,\text{unbiased}}+\theta_{t,\text{biased}}.$
+Specifically, we assume: (1) there exists a unique bias direction, $u$, such that $\theta$ can be decomposed into $\theta_{\text{unbiased}}$, which is orthogonal to $u$, and $\theta_{\text{biased}}$, where $|\theta_{\text{biased}}\cdot u|>0$; and (2) the extent of bias in the model is measured by $|\theta_{\text{biased}}\cdot u|$.
+During gradient-based optimization, the update rule is:
+$\theta_{t+1}=\theta_{t}-\eta\nabla_{\theta}\mathcal{L}_{\text{ft}}(\theta_{t}),$
+where $\eta$ is the learning rate, and $\mathcal{L}_{\text{ft}}$ denotes the fine-tuning loss function. Substituting the decomposition of $\theta_{t}$ and taking the projection, we have:
+$\theta_{t+1}=\theta_{t,\text{unbiased}}+\theta_{t,\text{biased}}-\eta\left(\frac{\mathbf{\theta_{t,\text{biased}}}}{\|\mathbf{\theta_{t,\text{biased}}}\|}\right)c_{t}$
+where $c_{t}$ is the bias projection coefficient, measuring the projection of the gradient onto the normalized biased component of the parameters:
+$c_{t}=\left(\frac{\theta_{t,\text{biased}}}{\|\theta_{t,\text{biased}}\|}\right)^{\top}\nabla_{\theta}\mathcal{L}_{\text{ft}}(\theta_{t}).$ (2)
+If $c_{t}<0$, the gradient update will reinforce the biased component, leading to bias amplification, i.e.
+$\Delta|\theta_{\text{biased}}|>0$.
+This occurs because the gradient descent step moves the parameters further in the direction of the existing bias.
+Another cause is sampling error, akin to statistical approximation error Shumailov et al.
+(2024).
+If the model has a pre-existing bias, it inherently assigns higher probabilities to tokens that produce biased outputs.
+Consequently, during synthetic data generation, unbiased tokens—and thus unbiased samples—are more likely to be lost at each resampling step with a finite sample, though this error vanishes as the sample size approaches infinity.
+This overrepresents biased patterns in the synthetic data, surpassing the model’s original bias and true next-token probabilities.
+Sampling error thus complements bias projection by further activating biased neurons in response to the skewed dataset.
+By definition, bias projection is a sufficient condition for bias amplification, while sampling error serves as a complementary factor.
+However, sampling error is a sufficient condition for model collapse to occur with nonzero probability Shumailov et al.
+(2024).
+This distinction might explain why bias amplification can occur without model collapse.
+I.2 Statistical Simulation
+To simulate a controlled setting without sampling error, we consider a statistical estimation cycle using WMLE with a large sample size of each resampling step.
+Specifically, we generate a pre-training dataset $\mathcal{D}_{\text{pre}}$ with 100,000 samples from a Beta$(3,2)$ distribution, representing a biased pre-training dataset.
+Using maximum likelihood estimation (MLE), we estimate its probability density function, yielding the pre-trained model $f_{\text{pre}}$.
+Next, we fine-tune $f_{\text{pre}}$ to approximate a different distribution, Beta(2,2)22(2,2)( 2 , 2 ).
+We generate 100,000 samples from this distribution, denoted as $D_{\text{real}}$, which serves as the initial fine-tuning dataset.
+In the first round, we apply weighted maximum likelihood estimation (WMLE) using weights derived from $f_{\text{pre}}$, which encode the pre-existing bias of the pre-trained model.
+This weighting captures the influence of the pre-trained model’s parameters on subsequent training.
+This produces the fine-tuned model $f_{\text{0}}$.
+We then generate a synthetic dataset $D_{\text{0}}$ of the same size using $f_{\text{0}}$, initiating the iterative fine-tuning loop.
+In each subsequent round, WMLE is applied using $D_{\text{k}}$ with weights from $f_{\text{k}}$, resulting in $f_{\text{k+1}}$.
+This process is repeated iteratively, producing models $f_{\text{1}}$ through $f_{\text{10}}$.
+Figure 10 shows the estimated distributions gradually shift toward the mean of the biased pre-training dataset at $x=0.6$, becoming progressively more peaked over generations.
+This occurs despite further training on samples drawn from Beta(2,2)22(2,2)( 2 , 2 ) and synthetic data generated from successive models.
+The distortion arises because the fine-tuning process disproportionately emphasizes regions where the pre-trained distribution assigns higher probability, leading to biased learning.
+Figure 10: Weighted Maximum Likelihood Estimation over 10 generations.
+For comparison, Figure 11 presents the results using standard MLE without weighting.
+In this case, the estimated distributions remain stable across generations, accurately representing the Beta(2,2)22(2,2)( 2 , 2 ) distribution.
+Figure 11: Maximum Likelihood Estimation over 10 generations.